### v1

# Evaluation of Dataframe Libraries for Data Preparation on a Single Machine

Abstract
Data preparation is a trial-and-error process that typically involves countless iterations over the data to define the best pipeline of operators for a given task. With tabular data, practitioners often perform that burdensome activity on local machines by writing ad hoc scripts with libraries based on the Pandas dataframe API and testing them on samples of the entire dataset—the faster the library, the less idle time its users have.
In this paper, we evaluate the most popular Python dataframe libraries in general data preparation use cases to assess how they perform on a single machine. To do so, we employ 4 real-world datasets and pipelines with distinct characteristics, covering a variety of scenarios. The insights gained with this experimentation are useful to data scientists who need to choose which of the dataframe libraries best suits their data preparation task at hand. In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when RAM is limited and there is no need to complete compatibility with Pandas API, Polars is the go-to choice thanks to its resource and query optimization; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multi-thread execution and a query optimizer) and Vaex (exploiting a columnar data format) are the best options.
Index Terms:
data preparation, dataframe, scalability
I Introduction
Companies and organizations significantly depend on their data to drive informed decisions, such as business strategy definition, supply chain management, etc. Thus, guaranteeing high data quality is fundamental for ensuring the reliability of analysis and avoiding undesired additional costs [1]. Oftentimes data is heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as the presence of missing or duplicate values [2]. Due to that, data scientists dedicate considerable time and resources to perform data preparation [3]. Typically, this fundamental process involves multiple operations called preparators [4], combined into a pipeline [5], which aim at exploring [6], cleaning [7], and transforming raw data into curated datasets [8, 9, 10].
It is said that data scientists spend up to 80% of their time on data preparation [11]. This also depends on the lack of support in the choice among many different available libraries and in the design of the pipeline that best suits both the dataset and the task at hand. The growing need for standardization and best practices has driven an increasing number of data practitioners to embrace the dataframe as the foundational data structure for constructing pipelines and developing libraries. At a high level, the dataframe is a two-dimensional data structure that consists of rows and columns [12]. Libraries for data preparation typically offer a flexible functional interface that allows to conveniently compose pipelines.
In particular, Pandas [13] is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators [12]. This library has become so popular in the data science ecosystem that even the popularity of Python has been attributed sometimes to its adoption [14]. Despite its success, Pandas notoriously presents some severe limits [15]. Firstly, it has not been designed to work with large datasets efficiently. In fact, Pandas operates in a single-threaded manner, lacks support for cluster deployment, and does not implement any memory optimization mechanism (the entire data is kept in main memory until the end of the pipeline execution). To overcome these limitations, several alternative libraries have been developed and are gaining popularity among data scientists. However, these libraries come with heterogeneous features (e.g., lazy evaluation, GPU support, etc.), making it hard for data scientists to navigate among them and choose the most suitable solution.
A major reason for this confusion is the lack of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario. Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons [16, 17, 18, 19, 20] (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature [21, 22, 23] either focus on orthogonal aspects or only cover small subsets of libraries and preparators. Similarly, relational benchmarking systems such as TPC-H [24] offer limited insights about data preparation workflows [22].
TABLE I: Features of the compared dataframe libraries.
Pandas PySpark Modin Polars CuDF Vaex DataTable
Multithreading
GPU acceleration
Resource optimization
Lazy evaluation
Deploy on cluster
[…]
Licence 3-Clause BSD Apache 2.0 Apache 2.0 MIT Apache 2.0 MIT Mozilla Public 2.0
[…]
Considered version 1.5.1-2.0.0 3.4.1 0.16.2 0.15.1 22.12.0 4.16.0 1.0.0
Our contribution
It is common for data scientists to perform countless iterations on samples of the datasets on their laptop/PC to define a proper data preparation pipeline for the task at hand. Hence, the faster the library employed for the pipeline execution, the less idle time they have. To support data scientists in the choice of the dataframe library that can speedup their workflow, we present the first extensive experimental comparison of the most popular alternative libraries to Pandas, namely PySpark, Modin (with Dask and Ray engines), Polars, CuDF, Vaex, and DataTable, including in the evaluation also Pandas itself and its evolution (Pandas2). For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library. Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries, and enables the execution in a containerized environment using Docker. This process is streamlined with the help of a convenient configuration file.
To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, collected from Kaggle [25] (the most popular platforms for data science challenges) and previously adopted in related work about data preparation [4] and dataframe libraries [21]. The choice of these datasets from Kaggle allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and that have been validated by the community.
We identify four essential data preparation stages in the pipelines: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC). Thus, we evaluate the performance of the different libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
For whom is this paper relevant
We focus only on the single-machine scenario. Our findings are useful to those data scientists that: (i) are employing Pandas in their workflow when devising data preparation pipelines on their laptop/pc; (ii) want to replace Pandas with a more efficient library that has the same interface. Thus, we evaluated scalability by measuring the performance of the dataframe libraries on three distinct machine configurations (laptop, workstation, and server), while increasing the size of the dataset. Hence, we provide the readers with several insights about their performance based on multiple factors, such as the size and the features of the dataset, the configuration of the underlying machine, and the type of applied preparators.
Outline
We describe in detail the dataframe data structure and the libraries to be compared in Section II, then we present our evaluation setup and the results in Section III and IV respectvely. We review the related literature in Section V, then we report the key takeaways in Section VI. In Section VII we draw the conclusions.
II Dataframes
A dataframe [12] is a data structure that organizes the data into a two-dimensional table-like structure, where columns represent the schema and rows represent the content. Each column has a specific data type, and the schema does not need to be declared in advance. A dataframe provides a wide set of operators acting on all data dimensions, enabling and simplifying various data preparation tasks such as data analysis, transformation, and cleaning. Beyond this variety of features, dataframes became popular for data analysis and manipulation thanks to their simple and flexible structure, which allows them to effectively handle different types of data. In fact, compared to using specific data preparation tools [4], dataframes offer greater simplicity and customization, enabling users to manipulate the data by writing Python code without having the effort of learning new tools.
Pandas [26] has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language [13]) and provides efficient and expressive data structures optimized to work with structured datasets. Pandas has established itself over time as the standard library for data manipulation and analysis. The main reasons behind its widespread popularity reside in its robust functionality and user-friendly design, making it the go-to choice for handling datasets in Python. Given its enormous diffusion, Pandas is a very mature library and the existence of a huge amount of documentation and courses make its learning by data scientists and practitioners extremely straightforward.
Nevertheless, Pandas notoriously presents several limitations [15], which become especially evident when dealing with large-scale datasets, as it lacks the optimizations required to process large amounts of data in an efficient way. For instance, Pandas does not support multiprocessing and parallel computing, and it is designed to work with in-memory datasets, lacking support for datasets that exceed memory limits. Further, it needs to materialize the intermediate result of every operation (this execution strategy is known as eager evaluation), exposing it to out-of-memory risks and preventing it from applying query optimization techniques.
While Pandas initially relied on NumPy [27] data structures for memory management, it moved to Apache Arrow [28] starting from Pandas2 [29]. Arrow is a language-independent columnar memory format for flat and hierarchical data, organized to optimize analytic operations on modern hardware, allowing Pandas2 to handle large datasets more efficiently.
To address the aforementioned limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing. In the following, we describe several of these popular Python libraries that can serve as replacements for Pandas. The selection of the libraries to include in our evaluation was based on three main characteristics they must have: (i) based on the dataframe data structure; (ii) compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the current reference standard, already aim at Pandas-compatibility for their API [30] or describe the parallelism between their API and the Pandas API [31]); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support. These libraries are: Pandas [26], PySpark [32], Modin [33], Polars [34], CuDF [35], Vaex [36], and DataTable [37]. We compare the selected libraries highlighting their strengths and weaknesses, to facilitate efficient data processing and provide support to data scientists and practitioners for efficiently handling their data preparation workflows.
TABLE II: Compatibility of dataframe libraries with Pandas API. (✓✓) fully matches Pandas interface; (✓) different interface; ($\circ$) missing from the API, but implemented by us to the best of our efforts.
Preparator Pandas API SparkPD SparkSQL Modin Polars CuDF Vaex DataTable
I/O load dataframe read_csv ✓✓ ✓✓ ✓✓ ✓✓
output dataframe to_csv ✓✓ ✓✓ ✓✓ ✓✓
EDA locate missing values (nulls) isna ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
locate outliers percentile ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
search by pattern str.contains ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
sort values sort ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get columns list columns ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get columns types dtypes ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get dataframe statistics describe ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
query columns query ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
DT cast columns types astype ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
delete columns drop ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
rename columns rename ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
pivot pivot_table ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
calculate column using expressions apply columnwise ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
join dataframes merge ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ $\circ$
one hot encoding get_dummies ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ $\circ$
categorical encoding cat.codes ✓✓ ✓✓ ✓✓ $\circ$
group dataframe groupby ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
DC change date & time format to_datetime ✓✓ ✓✓ $\circ$ ✓✓ $\circ$ $\circ$
delete empty and invalid rows dropna ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
set content case str.lower ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
normalize numeric values round ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
deduplicate rows drop_duplicates ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
fill empty cells fillna ✓✓ ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
replace values occurrences replace ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
edit & replace cell data apply rowise ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓
Dataframe Libraries
We identify five major features exploited by Pandas alternatives to achieve high efficiency:
  • Multithreading: using multithreading to speed up the execution of dataframe operations, making the most of the available hardware resources.
  • GPU acceleration: leveraging the parallel computing potential of graphics processing units to further improve the performance of dataframe operations.
  • Resource optimization: implementing strategies to efficiently manage memory, reducing the impact on the available resources, thus improving the overall performance.
  • Lazy evaluation: enhancing the overall efficiency by maintaining a logical plan of the operations, triggering their execution only when a specific output operation is invoked, to allow applying query optimization techniques.
  • Deploy on cluster: enabling the distribution of dataframe processing across a cluster of machines, leveraging parallel computing to enhance the overall performance (as stated in Section I, we aim to include distributed computing in our evaluation in future work).
Table I lists for each dataframe library which of these features it implements, along with other characteristics about its implementation and the version considered here.
PySpark [32] is the Python API for Apache Spark, which is primarily implemented in Scala. Spark provides the capability to work with distributed collections of Java or Python objects partitioned across a cluster, known as Relational Distributed Datasets (RDDs) [38]. PySpark dataframes are RDDs consisting of Row objects: they can be seen as relational tables, on which relational operators can be applied. Spark has been popularized by its capability to process large-scale datasets on a cluster of machines (represented within Spark as a SparkContext object). Yet, standalone mode (i.e., on a single machine) is not only supported, but is also surprisingly fast in multiple scenarios, as shown in the experimental evaluation of Section IV. PySpark dataframes support lazy evaluation, relying on a query optimizer called Catalyst [38]. Basically, Catalyst translates the pipeline of operators applied to a dataframe into a query tree of logical operators, which is then optimized to generate the final execution plan.
PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs for working with structured data, to efficiently read, write, transform, and analyze data using Python and SQL; (ii) Pandas on Spark (denoted as SparkPD and known as Koalas [39] before its inclusion in PySpark starting from Spark 3.2), which enables (by adding an index to the conventional Spark dataframe) the scaling of Pandas workloads by distributing them across multiple nodes without requiring modifications to the original Pandas code for most Pandas API functions ($\sim$80%).
Modin [33] is a Python library that provides a parallel alternative to Pandas [12]. Modin adopts the Pandas data format as the default storage layer, ensuring compatibility and usability, and parallelizes the execution of the operators across multiple cores [21]. Modin is designed to dynamically switch between different partition schemes: row-based, column-based, or block-based, depending on the operation. Each partition is then processed independently by the execution engine: Dask [40] or Ray [41].
Modin employs a set of 15 main operators, called core operators, to simplify Pandas functions and build its own Pandas-like API. When core operators cannot handle a function, Modin switches to the default to Pandas mode (incurring performance costs due to communication and Pandas single-threaded nature), then it reverts to a partitioned Modin dataframe after completion. Modin addresses flexible dataframe schemas by revising schema induction rules, allowing schema omission when not needed, optimizing queries, and minimizing overhead. The system implements opportunistic evaluation [42], enabling execution based on interactions. Incremental query construction is facilitated through intermediate dataframe results, prioritizing materialization of small results and employing predictive mechanisms. Modin captures and stores ordering information separately for on-demand ordering, and it explores key-value pairs for the efficient storage of sparse dataframes.
Since Modin can run using two different engines, Dask and Ray, we consider both in the evaluation (denoted as ModinD and ModinR, respectively). Dask [43] is a Python library for distributed computing, which allows working with large distributed Pandas dataframes. Dask uses dictionaries, tuples, and callables to represent task schedules. The scheduler can operate in different modes, providing flexibility to the execution. It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory. We considered the inclusion of Dask in our evaluation also as an independent library, but we found it to perform better in combination with Modin (Dask is not well-suited for a single machine), hence we do not report its results here. Furthermore, Dask covers approximately 55% of the Pandas API, while Modin around 90%.
Ray [44] is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Arrow as data format. Unlike Dask, Ray does not have built-in primitives for partitioned data. While the two engines serve different use cases, their primary goal is similar: to optimize resource usage by changing how data is stored and Python code is executed.
Polars [34] is a Python library written in Rust and built on top of arrow2 (the Rust implementation of Arrow format). The adoption of Arrow as underlying data structure confers efficient data processing capabilities to Polars, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
Polars does not use an index, but each row is indexed by its integer position in the table. Polars has developed its own Domain Specific Language (DSL) for transforming data, which aims to be very usable and ensure human readability even for complex queries. Its core components are Expressions and Contexts. Expressions facilitate concise and efficient data transformations, while Contexts categorize evaluations into three main types: (i) filtering, (ii) grouping/aggregation, (iii) selection. Polars optimizes queries through early filters and projection pushdown. Moreover, it supports both eager and lazy execution. The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can be executed in batches, lightening the load on memory and CPU, hence allowing processing bigger (and even larger-than-memory) datasets in a faster manner.
CuDF [35] is a component of the RAPIDS framework [45], developed by NVIDIA, which offers a Pandas-like API and enables to run general-purpose data science pipelines on GPUs. Written in C/C++ and CUDA, it harnesses the computational power of GPUs for accelerating data processing.
CuDF is built on top of Arrow and facilitates the interchange of tabular data across GPU processes. CuDF leverages parallelization to execute operations on different parts of columns simultaneously across GPU cores. This approach takes advantage of all available GPU cores, resulting in efficient and high-performance computations. Note that CuDF executes all operations on the GPU, but it does not provide any optimization strategy for the execution of the pipelines.
Vaex [36] is a Python library designed to handle extremely large tabular datasets (in particular astronomical catalogs), performing fast statistical analysis and visualization on this data. The key feature of Vaex is memory optimization, achieved through streaming algorithms, memory-mapped files, and a zero-copy policy, allowing to explore datasets exceeding the available memory, supported by its vaex-core extension module, written in C.
Vaex enables efficient column-wise operations by wrapping a series of NumPy arrays as columns. Only a small subset of NumPy functions benefits of lazy evaluation (e.g., exp, cos, arcsin), storing their results as computation instructions and computing them only when needed. This is achieved through virtual columns, expressions that can be added to datasets without incurring additional memory usage. Further, Vaex also supports multithreading for faster computation on multi-core CPUs. By default, Vaex avoids unnecessary data copying and loading into memory. It can process random data subsets and export data in random order for efficient resource utilization.
DataTable [37] is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns). While it shares similarities with Pandas, DataTable emphasizes speed and support for big data.
DataTable, which uses a native-C implementation for all data types, is developed to support column-oriented data storage, which optimizes data access and processing. DataTable enables memory-mapping of data on disk, allowing seamless processing of out-of-memory datasets, hence it is well-suited for large-scale data analysis tasks. DataTable incorporates multi-threaded data processing, leveraging all available cores for time-consuming operations, and offers efficient algorithms for sorting, grouping, and joining operations. To minimize unnecessary data copying, it adopts the copy-on-write technique for shared data, reducing memory overhead and improving overall performance.
TABLE III: Features of the selected datasets.
Athlete Loan Patrol Taxi CSV Size (GB) 0.03 1.6 6.7 10.9 # Rows (×$\times 10^{6}$) 0.2 2 27 77 # Columns 15 151 34 18 # Num - Str - Bool 5-10-0 113-38-0 5-27-2 15-3-0 % Null 9% 31% 22% 0% Str Len Range (1, 108) (1, 3988) (1, 2293) (1, 19)
III Bento
III-A Datasets
To evaluate the performance of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets that cover a variety of sizes, complexities, and characteristics, ensuring a comprehensive evaluation of the libraries: (i) Athlete [46], collecting 0.2 million records covering the competition results achieved by the athletes through 120 years of Olympic Games history; (ii) Loan [47], containing 2 million rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol [48], composed of 27 million records about 11 years of traffic stops by California law enforcement agencies; (iv) Taxi [49], containing 77 million records about taxi trips in New York City in 2015, hence also suitable for evaluating scalability. A summary of their main features is displayed in Table III.
Collected from Kaggle [25], all of these datasets were previously adopted in related work about data preparation and dataframe libraries. For instance, the three largest datasets are the ones selected by Petersohn et al. [21] to evaluate the performance of Modin, due to the variety of their features. Instead, Athlete was picked by Hameed and Naumann [4] to analyze how different preparators are supported by commercial data preparation tools.
(a) Athlete, EDA
(b) Loan, EDA
(c) Patrol, EDA
(d) Taxi, EDA
(e) Athlete, DT
(f) Loan, DT
(g) Patrol, DT
(h) Taxi, DT
(i) Athlete, DC
(j) Loan, DC
(k) Patrol, DC
(l) Taxi, DC
Figure 1: Average runtime for each stage of the data preparation pipeline (EDA, DT, DC) on different datasets.
III-B Data Preparation Pipelines
For each dataset, we selected from Kaggle the most voted notebook that proposed a non-trivial solution for the task at hand (i.e., Kaggle competitions) and extracted its data preparation pipeline—typically the first part of the notebook, preceding the definition of a machine learning model.
Overall, the pipelines employ a set of 27 preparators (reported in Table II) that can be clustered into four main stages:
  • Input/output (I/O): preparators that handle the input/output of data in various formats, such as databases, CSV or Parquet [50] files, Web APIs, etc.
  • Exploratory data analysis (EDA): preparators that support the exploratory analysis of the features of the data, for a better understanding of the data at hand and the detection of errors or anomalies.
  • Data transformation (DT): preparators that transform the data to make it more suitable for analysis, such as categorical encoding, join, aggregation, etc.
  • Data cleaning (DC): preparators to ensure the quality of the data at hand, thus the reliability of the results of its analysis (e.g., handling missing or erroneous data and correcting or removing outliers).
III-C Pipeline Specification
Bento is a Python framework that provides a general shared interface to design the data preparation pipeline, ignoring the differences in the APIs of the considered libraries. For preparator names, we adopted the convention proposed by Hameed and Naumann [4]. Each pipeline can be defined through a JSON file by specifying the sequence of preparators to employ. Our tool automatically deploys the pipeline for every specific library.
[…]
Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each stage (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline. Finally, our framework also supports the pipeline execution in Docker containers. Docker provides benefits like specifying core and memory usage and isolating Python environments to prevent package conflicts.
(a) Athlete
(b) Loan
Figure 2: Speedup over Pandas for each preparator on Athlete and Loan.
III-D Hardware and Software
All measurements are obtained as the average value over three runs on a machine equipped with an Intel Xeon Silver 4116 CPU (24 threads running at 2.1 GHz) and 196 GB of RAM, using Python 3.9, when not specified otherwise. The machine also features a NVIDIA Tesla T4 GPU with 16 GB of RAM and CUDA 11.2. The Ray engine was configured with default settings, resulting in 24 worker threads running. The Dask engine was configured comparably, leading to the creation of 6 workers and 24 threads for each execution. To ensure accuracy and avoid warm-up overhead, the assessment of execution time occurs once the system (e.g., JVM) has completed its warming-up process. To assess the performance of the libraries across machines with different hardware specifications, we simulated three distinct machine configurations using Docker, as detailed in Table IV. Finally, we evaluated scalability on incremental samples of Taxi and Patrol.
(a) Patrol
(b) Taxi
Figure 3: Speedup over Pandas for each preparator on Patrol and Taxi.
Figure 4: Average runtime for the apply preparator on Patrol and Taxi (only libraries not raising OoM issues).
IV Evaluation Results
In this section, we report and analyze the detailed results of our extensive experimental evaluation of the presented dataframe libraries. Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks. Thus, our comparison is designed to assess the performance of the dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
In particular, we organize our evaluation in multiple subsections, each of them is designed to answer one of the following research questions:
  1. Q1.
    What is the performance of the dataframe libraries in running data preparation pipelines (composed of the preparators applied most frequently by the Kaggle data science community) on datasets of different size and features? (Section IV-A)
  2. Q2.
    How does the optimizations introduced by lazy evaluations impact on the performance of the libraries that support it? (Section IV-B)
  3. Q3.
    How do the libraries scale by increasing the size of the datasets and varying the configuration of the machine used to run the data preparation pipeline (i.e., from laptop to server)? (Section IV-C)
IV-A Evaluation on Data Preparation Pipelines
Summary—For EDA, Polars is generally the best performer, up to 10 times faster than Pandas. For DT and DC, if a GPU is available, CuDF generally outperforms other libraries, but SparkSQL and Vaex achieve the best results on the largest dataset. Despite not supporting the Parquet format, DataTable is the fastest library for I/O.
For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 shows the runtime for the execution of three main stages of the data preparation pipeline: EDA, DT, and DC. The I/O stage is considered separately in Figures 6 and 6. Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at stage level when supported. Additionally, in Figures 2 and 3, we present the independent performance of each distinct preparator (i.e., we force the execution separately for each preparator), normalized with respect to Pandas (the de facto standard library for managing dataframes). For each preparator, the red line represents therefore Pandas as the baseline, while distinct markers denote the speedup for each library, which is defined as follows:
[…]
In the next subsections, we analyze in detail the results obtained for each of the main data preparation stages, considering both the single preparators and the entire stages. Note that not all preparators have the same impact on the pipeline, but their weight depends on the time required for their execution (as we point out when analyzing the related results).
Figure 5: Average runtime for reading CSV and Parquet files.
(a) Athlete, Read
(b) Loan, Read
(c) Patrol, Read
(d) Taxi, Read
(a) Athlete, Write
(b) Loan, Write
(c) Patrol, Write
(d) Taxi, Write
Figure 5: Average runtime for reading CSV and Parquet files.
Figure 6: Average runtime for writing CSV and Parquet files.
IV-A1 Explorative Data Analysis (EDA)
In this stage, Polars clearly stands out as the best performer on all datasets but Taxi, where ModinD outperforms other dataframe libraries (Figures 0(a)-0(d)). In this stage, four preparators alone account for 95% of the total execution time: locate missing values (isna), locate outliers (outlier), search by pattern (srchptn), and sort values (sort).
Note that Polars registers the best performance for two of them, locate missing values and locate outliers (Figures 2 and 3), where it is about ${10^{4}}$ times faster than Pandas and 0.3 times faster than the second best performer, DataTable. In fact, Polars can quickly locate missing values by exploiting the specific metadata stored by Arrow arrays [51]. Nevertheless, DataTable records a better performance in locating missing values on Patrol (Figure 2(a)), with a speedup of ${10^{5}}$ over Pandas and ${10^{3}}$ over Polars. As highlighted by Table III, Patrol is mostly composed of string attributes. For this type of data, Polars uses Arrow large_utf8 buffers, which are more expensive to filter than Python strings/chars, requiring therefore extra work.
For the other two most relevant preparators, in Figure 2(a) we can observe that Vaex excels in searching by pattern, being about 100 times faster than Pandas and 10 times faster than CuDF, which is the second best performer. Vaex tends to be very efficient in column-wise operations due to its internal columnar organization of the data. On the other hand, it is much less efficient when applying a row-wise preparator, for instance in locating missing values or outliers, where it is even about 100 times slower than Pandas. For sorting values, while CuDF is the fastest library on Athlete, PySpark (in both versions) shows better results as the datasets get larger. Note that Modin needs to convert the dataframe back and forth to Pandas for the sorting operation, hence it is generally the worst performer, up to 100 times slower than SparkSQL.
Finally, including preparators accounting for less than 5% of the total execution time, such as get columns list (gcols), get columns types (dtypes), get dataframe statistics (stats), and query columns (query), Modin outperforms other libraries in computing statistics exploiting its multithreading power, which even makes ModinD faster than Polars when running the entire stage on Taxi. Overall, Pandas maintains good performance throughout the stage, while Pandas2 shows slight but not significant improvements over its predecessor.
IV-A2 Data Trasformation (DT)
In this stage, CuDF emerges as the best performer on Athlete (Figure 0(e)) and Patrol (Figure 0(g)), while Polars and SparkSQL register the best performance (also thanks to lazy evaluation) on Loan (Figure 0(f)) and Taxi (Figure 0(h)), respectively.
All libraries show better memory management than Pandas, which goes out of memory even on Patrol, with the calculate column using expressions (apply) preparator. The comparison among the libraries that successfully complete their execution is shown in Figure 4, considering their absolute runtime (i.e., the lower the better). Except for Vaex, all libraries struggle with column-wise operations, where Vaex outperforms them by leveraging its internal columnar organization. Note that when running the entire stage (Figure 0(g)) Pandas has enough time to release memory, preventing out-of-memory issues. Pandas2 follows the trend of its predecessor, without any significant improvement.
Finally, the lazy execution of SparkSQL proves to be extremely efficient when running the entire data transformation stage on larger datasets. In fact, it almost reaches the performance of CuDF on Patrol (Figure 0(g)), with a speedup of ${10^{3}}$ over Pandas, and emerges as the best performer on Taxi (Figure 0(h)), where Pandas (along with Pandas2 and SparkPD), Modin, and Polars run into out-of-memory issues.
IV-A3 Data Cleaning (DC)
In this stage, the libraries that stand out the most are CuDF, and Vaex. In particular, CuDF consistently outperforms other libraries on Athlete and Loan, the two smallest datasets (Figures 0(i) and 0(j)). The pipelines designed for these two datasets include a preparator, deduplicate rows (dedup), accounting for 70% of the overall execution time for the data cleaning stage. Note that on Athlete CuDF is the only library performing better than Pandas (approximately 10 times faster), which on average maintains good performance, both in its basic version and as Pandas2. On Loan, it is possible to notice how Vaex equals or (even significantly) exceeds the performance of CuDF for multiple preparators, but the greater overall impact of the ones for which it is outperformed by CuDF, i.e., deduplicate rows, fill empty cells (fillna), and change date & time format (chdate), lead to a consistent difference in the overall execution time for this stage (Figure 0(j)).
However, results are different when moving to larger datasets (Figures 0(k) and 0(l)). CuDF excels on Patrol, where it is over 100 times faster than Pandas in both preparators, i.e., delete empty rows (dropna) and change date & time format (chdate), and slightly outperforming Vaex (in dropna) and CuDF. Nevertheless, its performance decrease on Taxi, where instead Vaex achieves the best results, being 10 times faster than it and more than 100 times faster than Pandas.
(a) Athlete
(b) Loan
(c) Patrol
(d) Taxi
Figure 7: Average runtime for the execution of the entire data preparation pipelines.
IV-A4 Input/Output (I/O)
Loading data represents a fundamental step for the success of any data-driven process. In particular, CSV has established itself as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion [52]. Nowadays, the diffusion of Arrow (whose adoption represents the main novelty of Pandas2 compared to its predecessor, as stated in Section II) is pushing towards an increasing adoption of the Parquet format [50], known for its efficient data compression. In particular, developers often load Parquet data into memory and convert it to Arrow [53], recognized as an ideal in-memory transport layer for such data [54].
Figures 6 and 6 show performance in read and write operations on both file formats, CSV and Parquet. DataTable, which maps the file into memory and navigates through it using pointers, records the best overall performance in reading files, but it does not support the Parquet format. Polars is the best performer in reading Parquet files due to its capability to read them directly into memory using Arrow. When dealing with small datasets like Athlete, Parquet and CSV exhibit almost the same performance (Figure 4(a)); in some cases (e.g., Vaex), the use of Parquet can even lead to worse performance, due to the need to decompress the data. When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with SparkSQL is over 10 times faster than reading CSV files (Figure 4(d)).
Write operations exhibit a wide range of performance, mainly due to the need of compressing data before saving. Parquet files come with additional metadata and encoding, which may introduce some overhead for the writing process. Parquet performs better than CSV for smaller datasets, but this advantage often diminishes as the dataset size increases. For instance, while on Patrol the Parquet format brings significant advantages to Pandas (outperforming CSV by a factor of 27) thanks to its efficient dataframe serialization, Modin and CuDF are much faster in writing CSV files (Figure 5(c)). Nevertheless, CSV format causes CuDF to run out of memory on the largest dataset (Figure 5(d)), while it succeeds in writing Parquet files.
TABLE IV: Specifications of each machine configuration.
Laptop Workstation Server # CPUs 8 16 24 # RAM (GB) 16 64 128 Dask (workers - threads) 4-8 4-16 6-24 Ray (workers) 8 16 24
IV-B Impact of Lazy Evaluation
SummaryCuDF significantly outperforms other libraries when running the entire pipeline, but SparkSQL excels on the largest dataset. Lazy evaluation brings benefits mostly on large datasets, improving performance by 70% on average (up to almost 130% in some cases) compared to its eager counterpart.
Figure 7 reports for each library the runtime for the execution of the entire data preparation pipeline. For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL and Polars), we also point out the difference compared to eager evaluation. Considering the overall performance on the entire pipeline, CuDF consistently excels, by effectively leveraging on the power of GPU. Pandas performs well on smaller datasets, but struggles on larger ones, running out of memory even on Patrol. This is due to its eager approach (which requires to calculate intermediate results), causing higher memory consumption [55]. Additionally, Vaex and DataTable perform well on Patrol thanks to parallel processing, which allows them to efficiently exploit available CPU cores and manage many string columns.
Lazy evaluation brings significant benefits on large datasets, such as Patrol (Figure 6(c)). SparkPD takes advantage of this strategy with Pandas-specific optimizations (e.g., vectorized operations and memory efficiency), improving performance by 62%. Further, the good performance of Polars can be attributed to its query optimization through streaming processing, early filters, and projection pushdown, resulting in a remarkable 126% performance improvement over its eager counterpart (Figure 6(c)). Finally, it can be noticed how long logical plans can limit the optimization introduced by lazy evaluation. Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement on Patrol limited to 31% and performance on Taxi almost unchanged.
(a) Laptop
(b) Workstation
(c) Server
Figure 8: Average runtime for running the entire pipeline on incremental samples of Taxi.
IV-C Scalability
SummarySparkSQL is the best library for scaling to large datasets on a single machine. For both Patrol and Taxi datasets, SparkSQL is the only one able to execute the entire pipelines with the minimum configuration. CuDF follows closely but requires a GPU.
In this section, we assess how dataframe libraries scale by varying the size of the dataset and the configuration of the underlying machine. In particular, we use three Docker containers, whose CPU and RAM are incrementally increased as described in Table IV to reproduce typical configurations for laptops, workstations, and servers. Figure 8 show the performance of the libraries for running the entire pipeline on incremental samples of Taxi. Note that Pandas and Modin represent both their variants, since they show the same performance. Overall, SparkSQL stands out as the best performer and is the only one capable of handling the entire dataset on a laptop configuration (Figure 7(a)).
Moving to a workstation configuration (Figure 7(b)), even CuDF and Vaex can complete the pipeline, with the latter providing similar performance to SparkSQL. Finally, DataTable manages to complete the pipeline only on a server configuration (Figure 7(c)), while Pandas, SparkPD, Modin, and Polars all fail much earlier. In particular, Polars shows a good trend (similar to CuDF without the need for a GPU), but already faces memory issues to handle 15% of the dataset. Pandas and SparkPD exhibit the poorest performance instead, followed by Modin.
Table V outlines instead the minimum configuration required by each specific library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi. On both datasets, as previously stated, only SparkSQL can complete the entire pipeline with the minimum configuration. Following closely is CuDF, which can handle the entire Patrol dataset with a laptop configuration, but needs a GPU, making it more expensive. Further, Vaex stands out on Patrol as one of the libraries that require the least resources, followed closely by SparkPD. ModinR registers a performance on Taxi similar to its Dask counterpart, but scales much better on Patrol. In fact, while Dask engine uses a centralized scheduler to share work across multiple cores, Ray uses distributed bottom-up scheduling. Indeed, Modin can run on top of Dask, but it was originally built to work with Ray, and such integration is still more mature. DataTable maintains a relatively consistent performance on both datasets. Instead, despite excelling in other experiments, scalability appears to be a significant weakness for Polars. Finally, Pandas continues to show severe limitations.
TABLE V: Minimum machine configuration for running the entire pipeline on incremental dataset samples. Laptop, Workstation, Server. See Table IV for the detailed configurations.
Patrol Taxi
% Sample 1% 5% 15% 25% 50% 75% 100% 1% 5% 15% 25% 50% 75% 100%
# Rows (M) 0.2 1.2 4.5 6.7 13.5 20.2 27 0.7 3.8 11.5 19.2 38.8 57.7 77
Pandas LP LP WS SV SV LP WS SV
SparkPD LP LP LP LP LP WS WS WS WS WS WS
SparkSQL LP LP LP LP LP LP LP LP LP LP LP LP LP LP
ModinD LP LP WS WS WS SV SV LP LP WS WS SV
ModinR LP LP LP LP WS WS SV LP LP WS WS SV
Polars LP LP LP WS WS WS SV LP LP WS
cuDF LP LP LP LP LP LP LP LP LP LP WS WS WS WS
Vaex LP LP LP LP LP WS WS LP LP LP WS WS WS WS
Datatable LP LP LP WS SV SV SV LP LP LP WS WS WS SV
V Related Work
To the best of our knowledge, our paper presents the first rigorous and extensive experimental comparison of existing dataframe libraries on data preparation tasks. In fact, public wisdom about such libraries is mostly scattered across several not peer-reviewed sources [16, 17, 18, 19, 20], which generally compare the performance of few libraries (e.g., Pandas vs PySpark vs Polars) on a handful of operations. The major claims by such analysis are mostly confirmed by our evaluation; nevertheless, despite often providing some useful insights, they only offer a partial and very fragmented knowledge, far from a complete and detailed overview covering all libraries and operations.
In literature, papers presenting new dataframe libraries or dataframe-based tools usually include a comparison against some dataframe libraries. It is the case for the AFrame data analysis package [56] (later extended to PolyFrame [55]), designed to efficiently manage large-scale dataframes using SQL++ queries. AFrame is evaluated against Pandas, Spark, and Pandas on Ray (i.e., ModinR) through a micro-benchmark using the synthetic Wisconsin benchmark dataset [57] to assess acceleration and scalability across distributed environments. Similarly, Petersohn et al. [21] evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas. However, the mentioned comparisons share significant limitations, since data preparation is not the main topic of the analysis and only a relatively small set of operations is considered.
Further, dataframe libraries are sometimes compared in data science benchmarks. For instance, TPC-H [58], a popular decision support benchmark including a suite of business-oriented ad hoc queries and concurrent data modifications, is used by Polars [24] to compare against Pandas, PySpark, Dask, Vaex, and DuckDB [59]. Similarly, H2O (creator of DataTable[60] uses a set of queries to asses the scalability of group by and join operations on popular database and dataframe-like systems. Nevertheless, such relational benchmarking systems notoriously offer very limited insights about real-world workflows executed on dataframes [22].
Sanzu [61] is instead designed to evaluate five popular data science frameworks and systems (R, Anaconda [62], Dask, PostgreSQL [63], and PySpark) on data processing and analysis tasks. Sanzu is composed of a micro-benchmark, to test basic operations in isolation on synthetic datasets of different sizes, and a macro-benchmark, to evaluate series of operations representing concrete data science scenarios on real-world datasets. FuzzyData [22] is a workflow generation system that allows to compare dataframe-based APIs on worflows composed of a small subset of operations. Finally, Shanbhag et al. [23] measure the energy consumption of 21 dataframe processing tasks on Pandas, Vaex, and Dask.
VI Key Takeaways
Our evaluation confirmed the several known limitations of Pandas. However, despite the promises of the many libraries designed to overtake it, there is no silver bullet for data preparation of tabular data with dataframes on a single machine. We can help data practitioners to decide which library to use by answering the following questions.
VI-A Does the dataset fit in the GPU memory?
If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API. However, CuDF lacks of an optimizer, thus libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations [20].
VI-B No complete pipeline execution?
If a user is only interested in an individual stage (I/O, EDA, data transformation, or data cleaning), DataTable stands out as the fastest library for I/O operations—for other tasks it might not be that mature, with many preparators to implement manually. Polars emerges as the best performer for EDA. However, for Data Transformation and Data Cleaning, SparkSQL and Vaex show the best results.
VI-C What is the size of the datasets?
With small datasets ($<500k$ rows), all of the libraries show similar average runtimes. Even if Polars is often recommended for such a scenario, Pandas (especially Pandas2) appears to be the most reasonable choice instead, since their performance does not differ substantially, at least not enough to learn a new API and replace entire parts of the pipeline [18]. In general, if memory and CPU usage are limited, Polars has the best overall performance.
For datasets of medium size (from 2 to 19 million rows) with dense numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed. On the other hand, for sparse datasets with many null values (e.g., Loan), Polars and SparkPD show good performance. Finally, for datasets with many string columns (e.g., 25% Patrol), SparkSQL appears to be the best choice.
For larger datasets (over 20 million rows), SparkSQL stands out as the best option [64]. Furthermore, when the pipeline requires dealing with heavy-duty column operations, particularly in datasets that have a multitude of string columns and a good amount of null values (e.g., Patrol), Vaex seems to be the best move—and there is no need to set up any environment, as required for PySpark.
VII Conclusion
We presented an experimental comparison of the most used dataframe-based libraries to help practitioners in selecting the best suitable to run their data preparation tasks on a single machine. To guarantee reasonably comparable results and increase usability, we developed Bento, a general framework for assessing library performance on four major data preparation stages: Input/Output (I/O), Exploratory Data Analysis (EDA), Data Transformation (DT), and Data Cleaning (DC). Our contribution includes a thorough analysis of performance of the most used dataframe libraries on heterogeneous datasets, that ranges in size, complexity, and characteristics and the relative most voted pipeline from Kaggle.com. The results of our experiments show that there is no clear winner among the libraries, thus we summarized useful takeaways for supporting the selection of the right library for a task and dataset at hand. As future work, we are planning to compare the libraries in a distributed setting and to study how machine learning can support users in suggesting optimal library combinations based on datasets, preparators, and previous executions.
Acknowledgment
We extend our appreciation to the master’s students of the 2021/2022 academic year in the Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions to testing the libraries. This work was partially supported by the Department of Engineering “Enzo Ferrari” within the FARD-2022 (FAR2022DIP_DIEF - Prot. n. 2985 del 27/07/2022) and FARD-2023 (FAR_DIP_2023_DIEF-SIMONINI) projects.

### v2

@@
-Evaluation of Dataframe Libraries for Data Preparation on a Single Machine
-Abstract
+and Reggio Emilia, Italy
+Abstract.
 Data preparation is a trial-and-error process that typically involves countless iterations over the data to define the best pipeline of operators for a given task.
@@
 In this paper, we evaluate the most popular Python dataframe libraries in general data preparation use cases to assess how they perform on a single machine.
-To do so, we employ 4 real-world datasets and pipelines with distinct characteristics, covering a variety of scenarios.
+To do so, we employ 4 real-world datasets and pipelines with heterogeneous features, covering a variety of scenarios.
 The insights gained with this experimentation are useful to data scientists who need to choose which of the dataframe libraries best suits their data preparation task at hand.
-In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when RAM is limited and there is no need to complete compatibility with Pandas API, Polars is the go-to choice thanks to its resource and query optimization; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multi-thread execution and a query optimizer) and Vaex (exploiting a columnar data format) are the best options.
-Index Terms:
-data preparation, dataframe, scalability
-I Introduction
+In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when data fits in RAM and there is no need for complete compatibility with Pandas API, Polars is the go-to choice thanks to its in-memory execution and query optimizations; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multithread execution and a query optimizer) proves to be the best options.
+copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 28th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06
+1. Introduction
 Companies and organizations significantly depend on their data to drive informed decisions, such as business strategy definition, supply chain management, etc.
-Thus, guaranteeing high data quality is fundamental for ensuring the reliability of analysis and avoiding undesired additional costs [1].
-Oftentimes data is heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as the presence of missing or duplicate values [2].
-Due to that, data scientists dedicate considerable time and resources to perform data preparation [3].
-Typically, this fundamental process involves multiple operations called preparators [4], combined into a pipeline [5], which aim at exploring [6], cleaning [7], and transforming raw data into curated datasets [8, 9, 10].
-It is said that data scientists spend up to 80% of their time on data preparation [11].
+Thus, guaranteeing high data quality is fundamental to ensure the reliability of analysis and avoid undesired additional costs (Haug et al., 2011).
+Oftentimes data is heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as missing or duplicate values (Fan, 2015).
+Due to that, data scientists dedicate considerable time and resources to perform data preparation (Fernandes et al., 2023).
+Typically, this fundamental process involves multiple operations called preparators (Hameed and Naumann, 2020), combined into a pipeline (Furche et al., 2016), aimed at exploring (Lissandrini et al., 2018), cleaning (Ilyas and Chu, 2019), and transforming raw data into curated datasets (Terrizzano et al., 2015; Rezig et al., 2019; Geerts et al., 2020).
+It is said that data scientists spend up to 80% of their time on data preparation (Hellerstein et al., 2018).
 This also depends on the lack of support in the choice among many different available libraries and in the design of the pipeline that best suits both the dataset and the task at hand.
 The growing need for standardization and best practices has driven an increasing number of data practitioners to embrace the dataframe as the foundational data structure for constructing pipelines and developing libraries.
-At a high level, the dataframe is a two-dimensional data structure that consists of rows and columns [12].
+At a high level, the dataframe is a two-dimensional data structure that consists of rows and columns (Petersohn et al., 2020).
 Libraries for data preparation typically offer a flexible functional interface that allows to conveniently compose pipelines.
-In particular, Pandas [13] is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators [12].
-This library has become so popular in the data science ecosystem that even the popularity of Python has been attributed sometimes to its adoption [14].
-Despite its success, Pandas notoriously presents some severe limits [15].
+In particular, Pandas (McKinney, 2010) is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators (Petersohn et al., 2020; Jindal et al., 2021; Psallidas et al., 2022; Lee et al., [n.
+d.]).
+This library has become so popular in the data science ecosystem that even the popularity of Python itself has been sometimes attributed to its wide adoption (Claburn, 2017).
+Despite its success, Pandas notoriously presents some severe limits (Emani et al., 2024; Kläbe and Hagedorn, 2021).
 Firstly, it has not been designed to work with large datasets efficiently.
@@
 A major reason for this confusion is the lack of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario.
-Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons [16, 17, 18, 19, 20] (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature [21, 22, 23] either focus on orthogonal aspects or only cover small subsets of libraries and preparators.
-Similarly, relational benchmarking systems such as TPC-H [24] offer limited insights about data preparation workflows [22].
-TABLE I: Features of the compared dataframe libraries.
+Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons (Schmitt, 2020; Alexander, 2023; Pinner, 2023) (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature (Petersohn et al., 2021; Rehman and Elmore, 2022; Shanbhag and Chimalakonda, 2023) either focus on orthogonal aspects or only cover small subsets of libraries and preparators.
+Our contribution
+It is common for data scientists to perform countless iterations on samples of the datasets on their laptop or PC to define a proper data preparation pipeline for the task at hand.
+Hence, the faster the employed library is, the less idle time they have.
+To support data scientists in the choice of the dataframe library that can speedup their workflow, we present the first extensive experimental comparison of Pandas and its most popular dataframe-based alternatives, namely PySpark, Modin (with both Dask and Ray engine), Polars, CuDF, Vaex, and DataTable.
+For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library.
+Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries and enables the execution in a containerized environment using Docker.
+This process is streamlined with the help of a convenient configuration file.
+To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, previously adopted in related literature on data preparation (Hameed and Naumann, 2020) and dataframe libraries (Petersohn et al., 2021).
+All these datasets are collected from Kaggle [^2] https://www.kaggle.com, the most popular platform for data science challenges.
+This allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and have been validated by the world’s largest data science community.
+We identify four essential data preparation stages in the pipelines: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC).
+Thus, we evaluate the performance of the libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
+For whom is this paper relevant
+Since it represents by far the most common situation for data scientists (Emani et al., 2024), we only focus on the single-machine scenario, aiming to explore the distributed scenario in a future extension [^3] Note that we decided to include libraries conceived for distributed execution, such as PySpark or Modin, since through multithreading and optimized execution plans they can improve the performance of dataframe operations even on a single machine..
+Our findings are useful to those data scientists that: (i) are employing Pandas in their workflow when devising data preparation pipelines on their laptop or PC; (ii) want to replace Pandas with a more efficient library that has the same interface.
+Thus, we evaluate scalability by measuring the performance of the dataframe libraries on three distinct machine configurations (laptop, workstation, and server), while increasing the size of the dataset.
+Hence, we provide the readers with several insights about their performance based on multiple factors, such as the size and the features of the dataset, the configuration of the underlying machine, and the type of preparators applied.
+Outline
+We describe in detail the dataframe data structure and the libraries to be compared in Section 2, then we present our evaluation setup and the results in Section 3 and 4, respectively.
+After reviewing the related literature in Section 5, we report the key takeaways in Section 6 and draw the conclusions in Section 7.
+2. Dataframes
+A dataframe (Petersohn et al., 2020) is a data structure that organizes the data into a two-dimensional table-like format, where columns represent the schema and rows represent the content.
+Each column has a specific data type, and the schema does not need to be declared in advance.
+A dataframe provides a wide set of operators acting on all data dimensions, enabling and simplifying various data preparation tasks such as data analysis, transformation, and cleaning.
+Beyond this variety of features, dataframes became popular thanks to their simple and flexible structure, which allows them to effectively handle different types of data.
+In fact, compared to specific data preparation tools (Hameed and Naumann, 2020), dataframes offer greater simplicity and customization, enabling users to manipulate the data by writing Python code without the effort of learning to use a new tool.
+Pandas  [^4] https://github.com/pandas-dev/pandas has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language (McKinney, 2010)) and provides efficient and expressive data structures optimized to work with structured datasets.
+Pandas has established itself over time as the standard library for data manipulation and analysis.
+The main reasons behind its widespread popularity reside in its robust functionality and user-friendly design, making it the go-to choice for handling datasets in Python.
+Given its enormous diffusion, Pandas is a very mature library and the existence of a huge amount of documentation and courses makes its learning extremely straightforward.
+Nevertheless, Pandas also presents several notorious limitations (McKinney, 2017), which become especially evident when dealing with large-scale datasets, as it lacks the optimizations required to process large amounts of data in an efficient way.
+For instance, Pandas does not support multiprocessing and parallel computing, and it is designed to work with in-memory datasets, lacking support for datasets that exceed memory limits.
+Further, it needs to materialize the intermediate result of every operation (this execution strategy is known as eager evaluation), exposing it to out-of-memory risks and preventing it from applying query optimization techniques.
+To address such limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing.
+The selection of the Python libraries to include in our evaluation was based on three main criteria, requiring them to: (i) be based on the dataframe data structure (excluding therefore SQL-based solutions such as DuckDB (Raasveldt and Mühleisen, 2019)); (ii) be compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the de facto standard, already aim at Pandas-compatibility for their API [^5] https://modin.readthedocs.io/en/latest or describe the parallelism between their API and the Pandas API [^6] https://datatable.readthedocs.io/en/latest/manual/comparison_with_pandas.html); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support.
+The selected libraries are: Pandas, PySpark, Modin, Polars, CuDF, Vaex, and DataTable.
+Table 1. Features of the compared dataframe libraries.
 Pandas PySpark Modin Polars CuDF Vaex DataTable
@@
 […]
-Considered version 1.5.1-2.0.0 3.4.1 0.16.2 0.15.1 22.12.0 4.16.0 1.0.0
-Our contribution
-It is common for data scientists to perform countless iterations on samples of the datasets on their laptop/PC to define a proper data preparation pipeline for the task at hand.
-Hence, the faster the library employed for the pipeline execution, the less idle time they have.
-To support data scientists in the choice of the dataframe library that can speedup their workflow, we present the first extensive experimental comparison of the most popular alternative libraries to Pandas, namely PySpark, Modin (with Dask and Ray engines), Polars, CuDF, Vaex, and DataTable, including in the evaluation also Pandas itself and its evolution (Pandas2).
-For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library.
-Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries, and enables the execution in a containerized environment using Docker.
-This process is streamlined with the help of a convenient configuration file.
-To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, collected from Kaggle [25] (the most popular platforms for data science challenges) and previously adopted in related work about data preparation [4] and dataframe libraries [21].
-The choice of these datasets from Kaggle allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and that have been validated by the community.
-We identify four essential data preparation stages in the pipelines: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC).
-Thus, we evaluate the performance of the different libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
-For whom is this paper relevant
-We focus only on the single-machine scenario.
-Our findings are useful to those data scientists that: (i) are employing Pandas in their workflow when devising data preparation pipelines on their laptop/pc; (ii) want to replace Pandas with a more efficient library that has the same interface.
-Thus, we evaluated scalability by measuring the performance of the dataframe libraries on three distinct machine configurations (laptop, workstation, and server), while increasing the size of the dataset.
-Hence, we provide the readers with several insights about their performance based on multiple factors, such as the size and the features of the dataset, the configuration of the underlying machine, and the type of applied preparators.
-Outline
-We describe in detail the dataframe data structure and the libraries to be compared in Section II, then we present our evaluation setup and the results in Section III and IV respectvely.
-We review the related literature in Section V, then we report the key takeaways in Section VI.
-In Section VII we draw the conclusions.
-II Dataframes
-A dataframe [12] is a data structure that organizes the data into a two-dimensional table-like structure, where columns represent the schema and rows represent the content.
-Each column has a specific data type, and the schema does not need to be declared in advance.
-A dataframe provides a wide set of operators acting on all data dimensions, enabling and simplifying various data preparation tasks such as data analysis, transformation, and cleaning.
-Beyond this variety of features, dataframes became popular for data analysis and manipulation thanks to their simple and flexible structure, which allows them to effectively handle different types of data.
-In fact, compared to using specific data preparation tools [4], dataframes offer greater simplicity and customization, enabling users to manipulate the data by writing Python code without having the effort of learning new tools.
-Pandas [26] has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language [13]) and provides efficient and expressive data structures optimized to work with structured datasets.
-Pandas has established itself over time as the standard library for data manipulation and analysis.
-The main reasons behind its widespread popularity reside in its robust functionality and user-friendly design, making it the go-to choice for handling datasets in Python.
-Given its enormous diffusion, Pandas is a very mature library and the existence of a huge amount of documentation and courses make its learning by data scientists and practitioners extremely straightforward.
-Nevertheless, Pandas notoriously presents several limitations [15], which become especially evident when dealing with large-scale datasets, as it lacks the optimizations required to process large amounts of data in an efficient way.
-For instance, Pandas does not support multiprocessing and parallel computing, and it is designed to work with in-memory datasets, lacking support for datasets that exceed memory limits.
-Further, it needs to materialize the intermediate result of every operation (this execution strategy is known as eager evaluation), exposing it to out-of-memory risks and preventing it from applying query optimization techniques.
-While Pandas initially relied on NumPy [27] data structures for memory management, it moved to Apache Arrow [28] starting from Pandas2 [29].
-Arrow is a language-independent columnar memory format for flat and hierarchical data, organized to optimize analytic operations on modern hardware, allowing Pandas2 to handle large datasets more efficiently.
-To address the aforementioned limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing.
-In the following, we describe several of these popular Python libraries that can serve as replacements for Pandas.
-The selection of the libraries to include in our evaluation was based on three main characteristics they must have: (i) based on the dataframe data structure; (ii) compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the current reference standard, already aim at Pandas-compatibility for their API [30] or describe the parallelism between their API and the Pandas API [31]); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support.
-These libraries are: Pandas [26], PySpark [32], Modin [33], Polars [34], CuDF [35], Vaex [36], and DataTable [37].
-We compare the selected libraries highlighting their strengths and weaknesses, to facilitate efficient data processing and provide support to data scientists and practitioners for efficiently handling their data preparation workflows.
-TABLE II: Compatibility of dataframe libraries with Pandas API.
+Considered version 2.2.1 3.5.1 0.29.0 0.20.23 24.04.01 4.17.0 1.1.0
+Dataframe Libraries
+We identify five major features exploited by Pandas alternatives to achieve high efficiency:
+  • Multithreading: using multithreading to speed up the execution of dataframe operations, making the most of the available hardware resources.
+  • GPU acceleration: leveraging the parallel computing potential of graphics processing units to further improve the performance of dataframe operations.
+  • Resource optimization: implementing strategies to efficiently manage memory, reducing the impact on the available resources, hence improving the overall performance.
+  • Lazy evaluation: maintaining a logical plan of the operations, triggering their execution only when a specific output operation is invoked, to apply query optimization techniques.
+  • Deploy on cluster: enabling the distribution of dataframe processing across a cluster of machines, leveraging parallel computing (as stated in Section 1, we aim to consider this dimension in future work).
+For each dataframe library, Table 1 lists which of these features it implements, along with information about its implementation and the version considered here.
+PySpark [^7] https://github.com/apache/spark is the Python API for Apache Spark, which is primarily implemented in Scala.
+Spark allows to work with distributed collections of Java or Python objects partitioned across a cluster, known as Relational Distributed Datasets (RDDs) (Armbrust et al., 2015).
+PySpark dataframes are RDDs consisting of Row objects: they can be seen as relational tables, on which relational operators can be applied.
+Further, they support lazy evaluation, relying on the Catalyst optimizer (Armbrust et al., 2015).
+While Spark owes its popularity to the capability of processing large-scale datasets on a cluster of machines, standalone (i.e., single-machine) mode is not only supported, but also surprisingly fast in multiple scenarios, as shown in Section 4.
+PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs for working with structured data; (ii) Pandas on Spark (denoted as SparkPD and previously known as Koalas [^8] https://github.com/databricks/koalas), which enables (by adding an index to the conventional Spark dataframe) to distribute Pandas workloads across multiple nodes without requiring modifications to the original Pandas code for most API functions ($\sim$80%).
+Table 2.
+Compatibility of dataframe libraries with Pandas API.
 (✓✓) fully matches Pandas interface; (✓) different interface; ($\circ$) missing from the API, but implemented by us to the best of our efforts.
@@
 edit & replace cell data apply rowise ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓
-Dataframe Libraries
-We identify five major features exploited by Pandas alternatives to achieve high efficiency:
-  • Multithreading: using multithreading to speed up the execution of dataframe operations, making the most of the available hardware resources.
-  • GPU acceleration: leveraging the parallel computing potential of graphics processing units to further improve the performance of dataframe operations.
-  • Resource optimization: implementing strategies to efficiently manage memory, reducing the impact on the available resources, thus improving the overall performance.
-  • Lazy evaluation: enhancing the overall efficiency by maintaining a logical plan of the operations, triggering their execution only when a specific output operation is invoked, to allow applying query optimization techniques.
-  • Deploy on cluster: enabling the distribution of dataframe processing across a cluster of machines, leveraging parallel computing to enhance the overall performance (as stated in Section I, we aim to include distributed computing in our evaluation in future work).
-Table I lists for each dataframe library which of these features it implements, along with other characteristics about its implementation and the version considered here.
-PySpark [32] is the Python API for Apache Spark, which is primarily implemented in Scala.
-Spark provides the capability to work with distributed collections of Java or Python objects partitioned across a cluster, known as Relational Distributed Datasets (RDDs) [38].
-PySpark dataframes are RDDs consisting of Row objects: they can be seen as relational tables, on which relational operators can be applied.
-Spark has been popularized by its capability to process large-scale datasets on a cluster of machines (represented within Spark as a SparkContext object).
-Yet, standalone mode (i.e., on a single machine) is not only supported, but is also surprisingly fast in multiple scenarios, as shown in the experimental evaluation of Section IV.
-PySpark dataframes support lazy evaluation, relying on a query optimizer called Catalyst [38].
-Basically, Catalyst translates the pipeline of operators applied to a dataframe into a query tree of logical operators, which is then optimized to generate the final execution plan.
-PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs for working with structured data, to efficiently read, write, transform, and analyze data using Python and SQL; (ii) Pandas on Spark (denoted as SparkPD and known as Koalas [39] before its inclusion in PySpark starting from Spark 3.2), which enables (by adding an index to the conventional Spark dataframe) the scaling of Pandas workloads by distributing them across multiple nodes without requiring modifications to the original Pandas code for most Pandas API functions ($\sim$80%).
-Modin [33] is a Python library that provides a parallel alternative to Pandas [12].
-Modin adopts the Pandas data format as the default storage layer, ensuring compatibility and usability, and parallelizes the execution of the operators across multiple cores [21].
-Modin is designed to dynamically switch between different partition schemes: row-based, column-based, or block-based, depending on the operation.
-Each partition is then processed independently by the execution engine: Dask [40] or Ray [41].
-Modin employs a set of 15 main operators, called core operators, to simplify Pandas functions and build its own Pandas-like API.
-When core operators cannot handle a function, Modin switches to the default to Pandas mode (incurring performance costs due to communication and Pandas single-threaded nature), then it reverts to a partitioned Modin dataframe after completion.
-Modin addresses flexible dataframe schemas by revising schema induction rules, allowing schema omission when not needed, optimizing queries, and minimizing overhead.
-The system implements opportunistic evaluation [42], enabling execution based on interactions.
-Incremental query construction is facilitated through intermediate dataframe results, prioritizing materialization of small results and employing predictive mechanisms.
-Modin captures and stores ordering information separately for on-demand ordering, and it explores key-value pairs for the efficient storage of sparse dataframes.
-Since Modin can run using two different engines, Dask and Ray, we consider both in the evaluation (denoted as ModinD and ModinR, respectively).
-Dask [43] is a Python library for distributed computing, which allows working with large distributed Pandas dataframes.
-Dask uses dictionaries, tuples, and callables to represent task schedules.
-The scheduler can operate in different modes, providing flexibility to the execution.
-It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory.
-We considered the inclusion of Dask in our evaluation also as an independent library, but we found it to perform better in combination with Modin (Dask is not well-suited for a single machine), hence we do not report its results here.
-Furthermore, Dask covers approximately 55% of the Pandas API, while Modin around 90%.
-Ray [44] is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Arrow as data format.
+Modin [^9] https://github.com/modin-project/modin is a Python library that provides a parallel alternative to Pandas (Petersohn et al., 2020, 2021).
+Modin adopts the Pandas data format as the default storage layer and employs a set of 15 core operators to simplify Pandas functions and build its own Pandas-like API.
+When core operators cannot handle a function, it switches to the default to Pandas mode (affecting performance due to communication costs and Pandas single-threaded nature), reverting to a partitioned Modin dataframe after completion.
+It also implements opportunistic evaluation (Xin et al., 2021), enabling execution based on interactions, and facilitates incremental query construction through intermediate dataframe results.
+Modin is designed to dynamically switch between different partition schemes (row-based, column-based, or block-based) depending on the operation.
+Each partition is then processed independently by the execution engine: Dask or Ray (we consider both solutions, denoted as ModinD and ModinR, respectively).
+Dask  [^10] https://github.com/dask/dask is a Python library for distributed computing, which allows working with large distributed Pandas dataframes (Rocklin, 2015).
+It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory, and its scheduler provides flexibility to the execution.
+While we also considered the inclusion of Dask as an independent library, we do not report its results since we found it to perform better in combination with Modin (as Dask is not well suitable for a single machine) and it only covers $\sim$55of the Pandas API (while Modin $\sim$90%).
+Ray  [^11] https://github.com/ray-project/ray is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Apache Arrow [^12] https://github.com/apache/arrow (a language-independent columnar memory format, recently adopted even by Pandas  [^13] https://pandas.pydata.org/docs/dev/whatsnew/v2.0.0.html) as data format (Moritz et al., 2018).
 Unlike Dask, Ray does not have built-in primitives for partitioned data.
-While the two engines serve different use cases, their primary goal is similar: to optimize resource usage by changing how data is stored and Python code is executed.
-Polars [34] is a Python library written in Rust and built on top of arrow2 (the Rust implementation of Arrow format).
+While the two engines serve different use cases, their primary goal is similar: optimizing resource usage by changing how data is stored and Python code is executed.
+Polars [^14] https://github.com/pola-rs/polars is a Python library written in Rust and built on top of arrow2 (the Rust implementation of Arrow format).
 The adoption of Arrow as underlying data structure confers efficient data processing capabilities to Polars, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
 Polars does not use an index, but each row is indexed by its integer position in the table.
-Polars has developed its own Domain Specific Language (DSL) for transforming data, which aims to be very usable and ensure human readability even for complex queries.
-Its core components are Expressions and Contexts.
-Expressions facilitate concise and efficient data transformations, while Contexts categorize evaluations into three main types: (i) filtering, (ii) grouping/aggregation, (iii) selection.
+Further, it has developed its own Domain Specific Language (DSL) for transforming data, whose core components are Expressions and Contexts.
+Expressions facilitate concise and efficient data transformations, while Contexts categorize evaluations into three main types: filtering, grouping/aggregation, and selection.
 Polars optimizes queries through early filters and projection pushdown.
 Moreover, it supports both eager and lazy execution.
-The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can be executed in batches, lightening the load on memory and CPU, hence allowing processing bigger (and even larger-than-memory) datasets in a faster manner.
-CuDF [35] is a component of the RAPIDS framework [45], developed by NVIDIA, which offers a Pandas-like API and enables to run general-purpose data science pipelines on GPUs.
-Written in C/C++ and CUDA, it harnesses the computational power of GPUs for accelerating data processing.
-CuDF is built on top of Arrow and facilitates the interchange of tabular data across GPU processes.
-CuDF leverages parallelization to execute operations on different parts of columns simultaneously across GPU cores.
-This approach takes advantage of all available GPU cores, resulting in efficient and high-performance computations.
-Note that CuDF executes all operations on the GPU, but it does not provide any optimization strategy for the execution of the pipelines.
-Vaex [36] is a Python library designed to handle extremely large tabular datasets (in particular astronomical catalogs), performing fast statistical analysis and visualization on this data.
-The key feature of Vaex is memory optimization, achieved through streaming algorithms, memory-mapped files, and a zero-copy policy, allowing to explore datasets exceeding the available memory, supported by its vaex-core extension module, written in C.
+The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can run in batches, lightening the load on memory and CPU, hence allowing to process bigger datasets (even larger than memory) in a faster manner.
+CuDF [^15] https://github.com/rapidsai/cudf is a component of the NVIDIA RAPIDS [^16] https://rapids.ai framework.
+Written in C/C++ and CUDA (hence only compatible with NVIDIA GPUs), it offers a Pandas-like API to run general-purpose data science pipelines on GPUs, leveraging their computational power for accelerating data processing.
+CuDF is built on top of Arrow and leverages parallelization to execute operations on different parts of columns simultaneously across all available GPU cores.
+Note that CuDF uses a single GPU, while Dask-CuDF can be used for multi-GPU parallel computing.
+Thus, it can perform efficient and high-performance computations, although it does not provide any optimization strategy for the execution of the pipelines.
+Vaex [^17] https://github.com/vaexio/vaex is a Python library [^18] Since the last commit on GitHub is dated about 10 months ago, the project might not be actively maintained at the moment.
+designed to handle extremely large tabular datasets (such as astronomical catalogs), producing fast statistical analysis and visualization.
+Supported by its vaex-core extension, written in C, it achieves memory optimization through streaming algorithms, memory-mapped files, and a zero-copy policy, allowing to explore datasets exceeding the available memory.
 Vaex enables efficient column-wise operations by wrapping a series of NumPy arrays as columns.
-Only a small subset of NumPy functions benefits of lazy evaluation (e.g., exp, cos, arcsin), storing their results as computation instructions and computing them only when needed.
+Only a small subset of NumPy functions benefits of lazy evaluation, storing their results as computation instructions and computing them only when needed.
 This is achieved through virtual columns, expressions that can be added to datasets without incurring additional memory usage.
-Further, Vaex also supports multithreading for faster computation on multi-core CPUs.
-By default, Vaex avoids unnecessary data copying and loading into memory.
-It can process random data subsets and export data in random order for efficient resource utilization.
-DataTable [37] is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns).
-While it shares similarities with Pandas, DataTable emphasizes speed and support for big data.
-DataTable, which uses a native-C implementation for all data types, is developed to support column-oriented data storage, which optimizes data access and processing.
-DataTable enables memory-mapping of data on disk, allowing seamless processing of out-of-memory datasets, hence it is well-suited for large-scale data analysis tasks.
-DataTable incorporates multi-threaded data processing, leveraging all available cores for time-consuming operations, and offers efficient algorithms for sorting, grouping, and joining operations.
-To minimize unnecessary data copying, it adopts the copy-on-write technique for shared data, reducing memory overhead and improving overall performance.
-TABLE III: Features of the selected datasets.
+Vaex can process random data subsets and export data in random order for efficient resource utilization.
+Further, it also supports multithreading for faster computation on multi-core CPUs.
+DataTable [^19] https://github.com/h2oai/datatable is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns).
+DataTable, which uses a native-C implementation for all data types, is developed to support column-oriented data storage, optimizing data access and processing.
+DataTable enables memory-mapping of data on disk, allowing the seamless processing of out-of-memory datasets.
+DataTable incorporates multithreaded data processing, leveraging all available cores for time-consuming operations, and offers efficient algorithms for sorting, grouping, and joining operations.
+To minimize unnecessary data copying, it adopts the copy-on-write technique for shared data, reducing memory overhead and improving the overall performance.
+3. Bento
+In this section, we provide all details about the configuration of Bento, our framework for evaluating different dataframe libraries on data preparation pipelines.
+Datasets
+For a comprehensive evaluation of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets covering a variety of sizes, complexities, and features: (i) Athlete  [^20] https://www.kaggle.com/datasets/heesoo37/120-years-of-olympic-history-athletes-and-results, collecting 0.2M records about the results achieved by the athletes through 120 years of Olympics; (ii) Loan  [^21] https://www.kaggle.com/datasets/wordsforthewise/lending-club, containing 2M rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol  [^22] https://www.kaggle.com/datasets/faressayah/stanford-open-policing-project, composed of 27M records about 11 years of traffic stops by California law enforcement agencies (Pierson et al., 2020); (iv) Taxi  [^23] https://www.kaggle.com/competitions/nyc-taxi-trip-duration, containing 77M records about taxi trips in New York City in 2015.
+A summary of their main features is displayed in Table 3.
+Collected from Kaggle, all of these datasets had previously been adopted in related work about data preparation and dataframe libraries.
+The three largest datasets were indeed used to evaluate the performance of Modin (Petersohn et al., 2021), due to the variety of their features.
+Instead, Athlete was exploited to analyze how different preparators are supported by commercial data preparation tools (Hameed and Naumann, 2020).
+Table 3. Features of the selected datasets.
 Athlete Loan Patrol Taxi CSV Size (GB) 0.03 1.6 6.7 10.9 # Rows (×$\times 10^{6}$) 0.2 2 27 77 # Columns 15 151 34 18 # Num - Str - Bool 5-10-0 113-38-0 5-27-2 15-3-0 % Null 9% 31% 22% 0% Str Len Range (1, 108) (1, 3988) (1, 2293) (1, 19)
-III Bento
-III-A Datasets
-To evaluate the performance of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets that cover a variety of sizes, complexities, and characteristics, ensuring a comprehensive evaluation of the libraries: (i) Athlete [46], collecting 0.2 million records covering the competition results achieved by the athletes through 120 years of Olympic Games history; (ii) Loan [47], containing 2 million rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol [48], composed of 27 million records about 11 years of traffic stops by California law enforcement agencies; (iv) Taxi [49], containing 77 million records about taxi trips in New York City in 2015, hence also suitable for evaluating scalability.
-A summary of their main features is displayed in Table III.
-Collected from Kaggle [25], all of these datasets were previously adopted in related work about data preparation and dataframe libraries.
-For instance, the three largest datasets are the ones selected by Petersohn et al.
-[21] to evaluate the performance of Modin, due to the variety of their features.
-Instead, Athlete was picked by Hameed and Naumann [4] to analyze how different preparators are supported by commercial data preparation tools.
+Data Preparation Pipelines
+For each dataset, we selected from Kaggle the most voted notebook that proposed a non-trivial solution for the task at hand (i.e., Kaggle competitions) and extracted its data preparation pipeline—typically the first part of the notebook, preceding the definition of a machine learning model.
+Overall, the pipelines employ a set of 27 preparators (reported in Table 2) that can be clustered into four main stages:
+  • Input/output (I/O): preparators to handle the input/output of data in various formats, such as databases, CSV or Parquet files, Web APIs, etc.
+  • Exploratory data analysis (EDA): preparators for supporting the exploratory analysis of data features, to better understand the data at hand and detect errors or anomalies.
+  • Data transformation (DT): preparators for transforming the data to make it more suitable for analysis, such as categorical encoding, join, aggregation, etc.
+  • Data cleaning (DC): preparators to improve the quality of the data, hence the reliability of the results of its analysis (e.g., handling missing or erroneous data and correcting or removing outliers).
+Pipeline Specification
+Bento is a Python framework that provides a general shared interface to design the data preparation pipeline, ignoring the differences in the APIs of the considered libraries.
+For preparator names, we adopted the convention proposed by Hameed and Naumann (Hameed and Naumann, 2020).
+Each pipeline can be defined through a JSON file by specifying the sequence of preparators to employ.
+Our tool automatically deploys the pipeline for every specific library.
+[…]
+Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each stage (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline.
+Finally, our framework also supports the pipeline execution in Docker containers, which allow to specify core and memory usage and isolate Python environments to prevent package conflicts.
+Hardware and Software
+All measurements are obtained as the average value over ten runs on a machine equipped with two AMD EPYC Rome 7402 CPUs (48 threads running at 2.8 GHz) and 512 GB of RAM, using Python 3.9, when not specified otherwise.
+The machine also features a NVIDIA A100 GPU with 40 GB of RAM and CUDA 12.1.
+The Ray engine was configured with default settings, resulting in 48 worker threads running.
+The Dask engine was configured comparably, leading to the creation of 8 workers and 48 threads for each execution.
+To ensure accuracy and avoid warm-up overhead, the assessment of execution time occurs once the system (e.g., JVM) has completed its warm-up process.
+To assess the performance of the libraries across machines with different hardware specifications, we simulated three distinct machine configurations, as detailed in Table 4.
+Finally, we evaluated scalability on incremental samples of Taxi and Patrol.
+4. Evaluation Results
+In this section, we report and analyze the results of the extensive experimental evaluation of the presented dataframe libraries.
+Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks.
+Thus, our comparison is designed to assess the performance of dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
 (a) Athlete, EDA
@@
 (l) Taxi, DC
-Figure 1: Average runtime for each stage of the data preparation pipeline (EDA, DT, DC) on different datasets.
-III-B Data Preparation Pipelines
-For each dataset, we selected from Kaggle the most voted notebook that proposed a non-trivial solution for the task at hand (i.e., Kaggle competitions) and extracted its data preparation pipeline—typically the first part of the notebook, preceding the definition of a machine learning model.
-Overall, the pipelines employ a set of 27 preparators (reported in Table II) that can be clustered into four main stages:
-  • Input/output (I/O): preparators that handle the input/output of data in various formats, such as databases, CSV or Parquet [50] files, Web APIs, etc.
-  • Exploratory data analysis (EDA): preparators that support the exploratory analysis of the features of the data, for a better understanding of the data at hand and the detection of errors or anomalies.
-  • Data transformation (DT): preparators that transform the data to make it more suitable for analysis, such as categorical encoding, join, aggregation, etc.
-  • Data cleaning (DC): preparators to ensure the quality of the data at hand, thus the reliability of the results of its analysis (e.g., handling missing or erroneous data and correcting or removing outliers).
-III-C Pipeline Specification
-Bento is a Python framework that provides a general shared interface to design the data preparation pipeline, ignoring the differences in the APIs of the considered libraries.
-For preparator names, we adopted the convention proposed by Hameed and Naumann [4].
-Each pipeline can be defined through a JSON file by specifying the sequence of preparators to employ.
-Our tool automatically deploys the pipeline for every specific library.
-[…]
-Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each stage (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline.
-Finally, our framework also supports the pipeline execution in Docker containers.
-Docker provides benefits like specifying core and memory usage and isolating Python environments to prevent package conflicts.
+Figure 1. Average runtime for each stage of the data preparation pipeline (EDA, DT, DC) on different datasets.
+In particular, we organize our evaluation in multiple subsections, each designed to answer one of the following research questions:
+  1. Q1.
+    What is the performance of the dataframe libraries in running data preparation pipelines on datasets of different size and features? (Section 4.1)
+  2. Q2.
+    How does lazy evaluation impact on the performance of the libraries that support it? (Section 4.2)
+  3. Q3.
+    How do libraries scale by varying the size of the dataset and the configuration of the underlying machine (i.e., from laptop to server)? (Section 4.3)
+  4. Q4.
+    How do libraries perform on the standard queries of the TPC-H benchmark? (Section 4.4)
+4.1. Evaluation on Data Preparation Pipelines
+{mdframed}
+Summary—For EDA, Polars is generally the best performer, up to 100 times faster than Pandas.
+For DT, if a GPU is available, CuDF generally outperforms other libraries.
+For DC, ModinR achieves notable results on the largest datasets.
+Finally, despite not supporting the Parquet format, DataTable is the fastest library for I/O.
+For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 shows the runtime for the execution of three main stages of the data preparation pipeline: EDA, DT, and DC.
+The I/O stage is considered separately in Figures 4 and 4.
+Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at stage level when supported.
+Additionally, in Figure 2 we present the independent performance of each distinct preparator (i.e., we force the execution separately for each preparator), normalized with respect to Pandas (the de facto standard library for managing dataframes).
+For each preparator, the red line represents therefore Pandas as the baseline, while distinct markers denote the speedup for each library, which is defined as follows:
+$speedup=\frac{Time{\langle\textrm{Pandas},\textit{{prep}}\rangle}}{Time{\langle\textit{{lib}},\textit{{prep}}\rangle}}$
+where $Time{\langle\textit{{lib}},\textit{{prep}}\rangle}$ is the time required by the library lib to run the preparator prep.
+Thus, a value above the red line denotes that the library outperforms Pandas; otherwise, it takes more time than Pandas to run the preparator.
+The logarithmic scale enables to easily assess the order of magnitude of the improvement (if any) of each library over Pandas.
+In the next subsections, we analyze in detail the results obtained for each of the main data preparation stages, considering both the single preparators and the entire stages.
+Note that not all preparators have the same impact on the pipeline, but their weight depends on the time required for their execution (as we point out when analyzing the related results).
 (a) Athlete
 (b) Loan
-Figure 2: Speedup over Pandas for each preparator on Athlete and Loan.
-III-D Hardware and Software
-All measurements are obtained as the average value over three runs on a machine equipped with an Intel Xeon Silver 4116 CPU (24 threads running at 2.1 GHz) and 196 GB of RAM, using Python 3.9, when not specified otherwise.
-The machine also features a NVIDIA Tesla T4 GPU with 16 GB of RAM and CUDA 11.2.
-The Ray engine was configured with default settings, resulting in 24 worker threads running.
-The Dask engine was configured comparably, leading to the creation of 6 workers and 24 threads for each execution.
-To ensure accuracy and avoid warm-up overhead, the assessment of execution time occurs once the system (e.g., JVM) has completed its warming-up process.
-To assess the performance of the libraries across machines with different hardware specifications, we simulated three distinct machine configurations using Docker, as detailed in Table IV.
-Finally, we evaluated scalability on incremental samples of Taxi and Patrol.
-(a) Patrol
-(b) Taxi
-Figure 3: Speedup over Pandas for each preparator on Patrol and Taxi.
-Figure 4: Average runtime for the apply preparator on Patrol and Taxi (only libraries not raising OoM issues).
-IV Evaluation Results
-In this section, we report and analyze the detailed results of our extensive experimental evaluation of the presented dataframe libraries.
-Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks.
-Thus, our comparison is designed to assess the performance of the dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
-In particular, we organize our evaluation in multiple subsections, each of them is designed to answer one of the following research questions:
-  1. Q1.
-    What is the performance of the dataframe libraries in running data preparation pipelines (composed of the preparators applied most frequently by the Kaggle data science community) on datasets of different size and features?
-(Section IV-A)
-  2. Q2.
-    How does the optimizations introduced by lazy evaluations impact on the performance of the libraries that support it? (Section IV-B)
-  3. Q3.
-    How do the libraries scale by increasing the size of the datasets and varying the configuration of the machine used to run the data preparation pipeline (i.e., from laptop to server)?
-(Section IV-C)
-IV-A Evaluation on Data Preparation Pipelines
-Summary—For EDA, Polars is generally the best performer, up to 10 times faster than Pandas.
-For DT and DC, if a GPU is available, CuDF generally outperforms other libraries, but SparkSQL and Vaex achieve the best results on the largest dataset.
-Despite not supporting the Parquet format, DataTable is the fastest library for I/O.
-For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 shows the runtime for the execution of three main stages of the data preparation pipeline: EDA, DT, and DC.
-The I/O stage is considered separately in Figures 6 and 6.
-Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at stage level when supported.
-Additionally, in Figures 2 and 3, we present the independent performance of each distinct preparator (i.e., we force the execution separately for each preparator), normalized with respect to Pandas (the de facto standard library for managing dataframes).
-For each preparator, the red line represents therefore Pandas as the baseline, while distinct markers denote the speedup for each library, which is defined as follows:
-[…]
-[…]
-In the next subsections, we analyze in detail the results obtained for each of the main data preparation stages, considering both the single preparators and the entire stages.
-Note that not all preparators have the same impact on the pipeline, but their weight depends on the time required for their execution (as we point out when analyzing the related results).
-Figure 5: Average runtime for reading CSV and Parquet files.
+(c) Patrol
+(d) Taxi
+Figure 2. Speedup over Pandas for each preparator on different datasets.
+Explorative Data Analysis (EDA)
+For EDA, Polars clearly stands out as the best performer on all datasets but Patrol, where CuDF outperforms other dataframe libraries (Figures 1(a)-1(d)).
+In this stage, four preparators alone account for 95% of the total execution time: locate missing values (isna), locate outliers (outlier), search by pattern (srchptn), and sort values (sort).
+Note that Polars generally registers the best performance for two of them, locate missing values and locate outliers (Figure 2), where it is up to ${10^{4}}$ times faster than Pandas and 0.6 times faster than the second best performer, DataTable.
+In fact, Polars can quickly locate missing values by exploiting the specific metadata stored by Arrow arrays [^24] https://pola-rs.github.io/polars/user-guide/expressions/null/.
+Nevertheless, DataTable is the fastest in locating missing values on Patrol (Figure 2(c)), with a speedup of ${10^{4}}$ over Pandas and ${10^{3}}$ over Polars, thanks to its implementation of binary-search-based filtering.
+As highlighted by Table 3, Patrol is mostly composed of string attributes.
+For this type of data, Polars uses Arrow large_utf8 buffers, which are more expensive to filter than Python strings/chars, requiring therefore extra work.
+For the other two most relevant preparators, in Figure 2(c) we can observe that Vaex excels in searching by pattern, being about ${10^{3}}$ times faster than Pandas and 10 times faster than the second best performer, CuDF.
+Vaex tends to be very efficient in column-wise operations due to its internal columnar organization of the data.
+On the other hand, it is much less efficient when applying a row-wise preparator, for instance in locating outliers, where it is even about 100 times slower than Pandas.
+Conversely, when locating missing values, Vaex shows relatively good performance on dense datasets such as Taxi and Patrol, though it is among the less efficient on sparse datasets like Loan, where its performance is does not provide improvements over Pandas (Figure 2(b)).
+For sorting values, CuDF and Polars are the fastest libraries overall.
+SparkSQL shows remarkable advantages over Pandas as the datasets grow larger, while SparkPD maintains similar performance to the baseline.
+Note that in Athlete, Modin converts by default the dataframe back and forth to Pandas for the sorting operation on multiple columns, hence in this case it is the worst performer, up to 100 times slower than Pandas.
+Finally, for the remaining preparators, get dataframe statistics (stats) and query columns (query), Modin outperforms other libraries on larger datasets exploiting its multithreading power, which even makes ModinD 10 times faster than Polars when running the entire stage on Taxi.
+Overall, Pandas maintains good performance throughout the stage.
 (a) Athlete, Read
@@
 (d) Taxi, Read
-(a) Athlete, Write
-(b) Loan, Write
-(c) Patrol, Write
-(d) Taxi, Write
-Figure 5: Average runtime for reading CSV and Parquet files.
-Figure 6: Average runtime for writing CSV and Parquet files.
-IV-A1 Explorative Data Analysis (EDA)
-In this stage, Polars clearly stands out as the best performer on all datasets but Taxi, where ModinD outperforms other dataframe libraries (Figures 0(a)-0(d)).
-In this stage, four preparators alone account for 95% of the total execution time: locate missing values (isna), locate outliers (outlier), search by pattern (srchptn), and sort values (sort).
-Note that Polars registers the best performance for two of them, locate missing values and locate outliers (Figures 2 and 3), where it is about ${10^{4}}$ times faster than Pandas and 0.3 times faster than the second best performer, DataTable.
-In fact, Polars can quickly locate missing values by exploiting the specific metadata stored by Arrow arrays [51].
-Nevertheless, DataTable records a better performance in locating missing values on Patrol (Figure 2(a)), with a speedup of ${10^{5}}$ over Pandas and ${10^{3}}$ over Polars.
-As highlighted by Table III, Patrol is mostly composed of string attributes.
-For this type of data, Polars uses Arrow large_utf8 buffers, which are more expensive to filter than Python strings/chars, requiring therefore extra work.
-For the other two most relevant preparators, in Figure 2(a) we can observe that Vaex excels in searching by pattern, being about 100 times faster than Pandas and 10 times faster than CuDF, which is the second best performer.
-Vaex tends to be very efficient in column-wise operations due to its internal columnar organization of the data.
-On the other hand, it is much less efficient when applying a row-wise preparator, for instance in locating missing values or outliers, where it is even about 100 times slower than Pandas.
-For sorting values, while CuDF is the fastest library on Athlete, PySpark (in both versions) shows better results as the datasets get larger.
-Note that Modin needs to convert the dataframe back and forth to Pandas for the sorting operation, hence it is generally the worst performer, up to 100 times slower than SparkSQL.
-Finally, including preparators accounting for less than 5% of the total execution time, such as get columns list (gcols), get columns types (dtypes), get dataframe statistics (stats), and query columns (query), Modin outperforms other libraries in computing statistics exploiting its multithreading power, which even makes ModinD faster than Polars when running the entire stage on Taxi.
-Overall, Pandas maintains good performance throughout the stage, while Pandas2 shows slight but not significant improvements over its predecessor.
-IV-A2 Data Trasformation (DT)
-In this stage, CuDF emerges as the best performer on Athlete (Figure 0(e)) and Patrol (Figure 0(g)), while Polars and SparkSQL register the best performance (also thanks to lazy evaluation) on Loan (Figure 0(f)) and Taxi (Figure 0(h)), respectively.
-All libraries show better memory management than Pandas, which goes out of memory even on Patrol, with the calculate column using expressions (apply) preparator.
-The comparison among the libraries that successfully complete their execution is shown in Figure 4, considering their absolute runtime (i.e., the lower the better).
-Except for Vaex, all libraries struggle with column-wise operations, where Vaex outperforms them by leveraging its internal columnar organization.
-Note that when running the entire stage (Figure 0(g)) Pandas has enough time to release memory, preventing out-of-memory issues.
-Pandas2 follows the trend of its predecessor, without any significant improvement.
-Finally, the lazy execution of SparkSQL proves to be extremely efficient when running the entire data transformation stage on larger datasets.
-In fact, it almost reaches the performance of CuDF on Patrol (Figure 0(g)), with a speedup of ${10^{3}}$ over Pandas, and emerges as the best performer on Taxi (Figure 0(h)), where Pandas (along with Pandas2 and SparkPD), Modin, and Polars run into out-of-memory issues.
-IV-A3 Data Cleaning (DC)
-In this stage, the libraries that stand out the most are CuDF, and Vaex.
-In particular, CuDF consistently outperforms other libraries on Athlete and Loan, the two smallest datasets (Figures 0(i) and 0(j)).
-The pipelines designed for these two datasets include a preparator, deduplicate rows (dedup), accounting for 70% of the overall execution time for the data cleaning stage.
-Note that on Athlete CuDF is the only library performing better than Pandas (approximately 10 times faster), which on average maintains good performance, both in its basic version and as Pandas2.
-On Loan, it is possible to notice how Vaex equals or (even significantly) exceeds the performance of CuDF for multiple preparators, but the greater overall impact of the ones for which it is outperformed by CuDF, i.e., deduplicate rows, fill empty cells (fillna), and change date & time format (chdate), lead to a consistent difference in the overall execution time for this stage (Figure 0(j)).
-However, results are different when moving to larger datasets (Figures 0(k) and 0(l)).
-CuDF excels on Patrol, where it is over 100 times faster than Pandas in both preparators, i.e., delete empty rows (dropna) and change date & time format (chdate), and slightly outperforming Vaex (in dropna) and CuDF.
-Nevertheless, its performance decrease on Taxi, where instead Vaex achieves the best results, being 10 times faster than it and more than 100 times faster than Pandas.
+Figure 3. Average runtime for reading CSV and Parquet files.
+(e) Athlete, Write
+(f) Loan, Write
+(g) Patrol, Write
+(h) Taxi, Write
+Figure 4. Average runtime for writing CSV and Parquet files.
+Data Trasformation (DT)
+In this stage, CuDF emerges as the best performer on Loan (Figure 1(f)) and Patrol (Figure 1(g)), while SparkSQL and Polars register the best performance (also thanks to lazy evaluation) on Athlete (Figure 1(e)) and Taxi (Figure 1(h)), respectively.
+The most time intensive preparators in this stage are categorical encoding (catenc), calculate new column (apply), and group dataframe (group).
+For catenc and group, CuDF always shows the best performance, being at least up to 100 times faster than the second best performer (Polars and ModinR, respectively).
+Once again, Vaex consistently outperforms the other libraries on column-wise operations (e.g., apply), being ${10^{3}}$ times faster than the second best performer CuDF (Figure 2(d)).
+On the other hand, this causes a drastic drop in its performance when moving to operations such as group, join, or pivot.
+As in the previous stage, SparkPD results as the worst performer on several preparators.
+This is mainly due to the latency introduced by translating Pandas operations into Spark execution plans.
+Since SparkSQL does not require this translation, it often achieves significantly better performance.
+When datasets scale up, libraries like ModinD and ModinR, designed for resource-intensive applications, exhibit superior performance due to their optimized handling of large-scale data compared to their performance on smaller datasets.
+Data Cleaning (DC)
+When moving to data cleaning operations, CuDF consistently outperforms other libraries on all datasets, except for Taxi (Figure 1(l)), where Vaex stands out as the best performer (with remarkable results achieved by Polars and Modin with both engines).
+The pipelines designed for the smallest datasets (Athlete and Loan) include a preparator, deduplicate rows (dedup), accounting for 70% of the overall execution time for the stage.
+Pandas register good results on these datasets, especially on Athlete, where only CuDF performs better (Figure 1(i)).
+On Loan, it is possible to notice how Vaex equals or (even significantly) exceeds the performance of CuDF for multiple preparators (Figure 2(b)), but the greater overall impact of the ones where it is outperformed by CuDF (especially the great difference in deduplicating rows) leads to a consistent difference in the overall execution time for this stage (Figure 1(j)).
+Results change when moving to larger datasets (Figures 1(k) and 1(l)).
+CuDF excels on Patrol, where it is over ${10^{3}}$ times faster than Pandas for both considered preparators, i.e., delete empty rows (dropna) and change date & time format (chdate), slightly outperforming ModinR.
+Nevertheless, it struggles on Taxi, where Vaex achieves the best results instead, followed closely by Modin (with both engines) and Polars.
 (a) Athlete
@@
 (d) Taxi
-Figure 7: Average runtime for the execution of the entire data preparation pipelines.
-IV-A4 Input/Output (I/O)
+Figure 5. Average runtime for the execution of the entire data preparation pipelines.
+Input/Output (I/O)
 Loading data represents a fundamental step for the success of any data-driven process.
-In particular, CSV has established itself as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion [52].
-Nowadays, the diffusion of Arrow (whose adoption represents the main novelty of Pandas2 compared to its predecessor, as stated in Section II) is pushing towards an increasing adoption of the Parquet format [50], known for its efficient data compression.
-In particular, developers often load Parquet data into memory and convert it to Arrow [53], recognized as an ideal in-memory transport layer for such data [54].
-Figures 6 and 6 show performance in read and write operations on both file formats, CSV and Parquet.
-DataTable, which maps the file into memory and navigates through it using pointers, records the best overall performance in reading files, but it does not support the Parquet format.
-Polars is the best performer in reading Parquet files due to its capability to read them directly into memory using Arrow.
-When dealing with small datasets like Athlete, Parquet and CSV exhibit almost the same performance (Figure 4(a)); in some cases (e.g., Vaex), the use of Parquet can even lead to worse performance, due to the need to decompress the data.
-When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with SparkSQL is over 10 times faster than reading CSV files (Figure 4(d)).
+In particular, the CSV format has established itself as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion (Vitagliano et al., 2023).
+Nowadays, the diffusion of Arrow is pushing towards an increasing adoption of the Parquet format [^25] https://parquet.apache.org, known for its efficient data compression.
+In particular, developers often load Parquet data into memory and convert it to Arrow (Liu et al., 2023), recognized as an ideal in-memory transport layer for such data [^26] https://arrow.apache.org/docs/python/parquet.html.
+Figures 4 and 4 show performance in read and write operations on both file formats, CSV and Parquet.
+DataTable, which maps the file into memory and navigates through it using pointers, records the best overall performance in reading CSV files, but it does not support the Parquet format.
+Polars is the best performer in reading Parquet files due to its capability to load them directly into memory using Arrow.
+When dealing with small datasets like Athlete, Parquet and CSV often exhibit almost the same performance, unlike Polars and SparkSQL (Figure 3(a)).
+In some cases (e.g., Vaex or CuDF), the use of Parquet can even lead to worse performance, due to the need for decompressing the data.
+When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with Polars is over 100 times faster than reading CSV files (Figure 3(d)).
 Write operations exhibit a wide range of performance, mainly due to the need of compressing data before saving.
 Parquet files come with additional metadata and encoding, which may introduce some overhead for the writing process.
-Parquet performs better than CSV for smaller datasets, but this advantage often diminishes as the dataset size increases.
-For instance, while on Patrol the Parquet format brings significant advantages to Pandas (outperforming CSV by a factor of 27) thanks to its efficient dataframe serialization, Modin and CuDF are much faster in writing CSV files (Figure 5(c)).
-Nevertheless, CSV format causes CuDF to run out of memory on the largest dataset (Figure 5(d)), while it succeeds in writing Parquet files.
-TABLE IV: Specifications of each machine configuration.
+Nevertheless, Parquet performs better than CSV across all datasets as it integrates various efficient compression and encoding approaches, while CSV is plain text and uncompressed (Liu et al., 2023).
+Table 4. Specifications of each machine configuration.
 Laptop Workstation Server # CPUs 8 16 24 # RAM (GB) 16 64 128 Dask (workers - threads) 4-8 4-16 6-24 Ray (workers) 8 16 24
-IV-B Impact of Lazy Evaluation
-SummaryCuDF significantly outperforms other libraries when running the entire pipeline, but SparkSQL excels on the largest dataset.
-Lazy evaluation brings benefits mostly on large datasets, improving performance by 70% on average (up to almost 130% in some cases) compared to its eager counterpart.
+4.2. Impact of Lazy Evaluation
+{mdframed}
+SummaryCuDF significantly outperforms other libraries when running the entire pipeline, while SparkSQL achieves remarkable results on all datasets.
+Lazy evaluation brings benefits mostly on large datasets, improving performance by 30% on average (up to almost 70% in some cases) compared to its eager counterpart.
 Figure 5 reports for each library the runtime for the execution of the entire data preparation pipeline.
-For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL and Polars), we also point out the difference compared to eager evaluation.
-Considering the overall performance on the entire pipeline, CuDF consistently excels, by effectively leveraging on the power of GPU.
-Pandas performs well on smaller datasets, but struggles on larger ones, running out of memory even on Patrol.
-This is due to its eager approach (which requires to calculate intermediate results), causing higher memory consumption [55].
-Additionally, Vaex and DataTable perform well on Patrol thanks to parallel processing, which allows them to efficiently exploit available CPU cores and manage many string columns.
-Lazy evaluation brings significant benefits on large datasets, such as Patrol (Figure 6(c)).
-SparkPD takes advantage of this strategy with Pandas-specific optimizations (e.g., vectorized operations and memory efficiency), improving performance by 62%.
-Further, the good performance of Polars can be attributed to its query optimization through streaming processing, early filters, and projection pushdown, resulting in a remarkable 126% performance improvement over its eager counterpart (Figure 6(c)).
-Finally, it can be noticed how long logical plans can limit the optimization introduced by lazy evaluation.
-Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement on Patrol limited to 31% and performance on Taxi almost unchanged.
+For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL, and Polars), we also point out the difference compared to eager evaluation.
+Considering the overall performance on the entire pipeline, CuDF consistently outperforms other libraries by effectively exploiting the power of GPU, with the only exception of the smallest dataset (Figure 5(a)).
+Pandas performs quite well on smaller datasets, but struggles on larger ones due to its eager approach (which requires computing intermediate results) causing higher memory consumption (Sinthong and Carey, 2021).
+Vaex and SparkSQL perform well on Patrol thanks to parallel processing, which allows them to efficiently exploit available CPU cores for handling many string columns.
+On one hand, it can be noticed how long logical plans can limit the optimizations introduced by lazy evaluation.
+Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement of 30% on Taxi and almost no changes in performance on Patrol.
+On the other hand, libraries like SparkPD leverage Pandas-specific optimizations (e.g., vectorized operations and memory efficiency) that translate into a 70% performance improvement.
+Further, Polars generally demonstrates good performance.
+In many cases, the eager API internally utilizes the lazy API before immediately collecting the result.
+Lazy execution in Polars leverages techniques such as streaming processing, early filtering, and projection pushdown, achieving a performance improvement of up to 15%.
+However, it is important to notice that these optimizations are not always available, hence in some cases lazy operations may perform comparably to their eager counterpart (Figures 5(b) and 5(d)).
 (a) Laptop
@@
 (c) Server
-Figure 8: Average runtime for running the entire pipeline on incremental samples of Taxi.
-IV-C Scalability
+Figure 6. Average runtime for running the entire pipeline on incremental samples of Taxi.
+4.3. Scalability
+{mdframed}
 SummarySparkSQL is the best library for scaling to large datasets on a single machine.
-For both Patrol and Taxi datasets, SparkSQL is the only one able to execute the entire pipelines with the minimum configuration.
-CuDF follows closely but requires a GPU.
+Indeed, for both Patrol and Taxi, it is the only one able to execute the entire pipelines with the laptop configuration.
+Polars follows closely, but requires a lot of resources in terms of RAM, exhibiting the worst scalability on machines with limited resources.
 In this section, we assess how dataframe libraries scale by varying the size of the dataset and the configuration of the underlying machine.
-In particular, we use three Docker containers, whose CPU and RAM are incrementally increased as described in Table IV to reproduce typical configurations for laptops, workstations, and servers.
-Figure 8 show the performance of the libraries for running the entire pipeline on incremental samples of Taxi.
-Note that Pandas and Modin represent both their variants, since they show the same performance.
-Overall, SparkSQL stands out as the best performer and is the only one capable of handling the entire dataset on a laptop configuration (Figure 7(a)).
-Moving to a workstation configuration (Figure 7(b)), even CuDF and Vaex can complete the pipeline, with the latter providing similar performance to SparkSQL.
-Finally, DataTable manages to complete the pipeline only on a server configuration (Figure 7(c)), while Pandas, SparkPD, Modin, and Polars all fail much earlier.
-In particular, Polars shows a good trend (similar to CuDF without the need for a GPU), but already faces memory issues to handle 15% of the dataset.
-Pandas and SparkPD exhibit the poorest performance instead, followed by Modin.
-Table V outlines instead the minimum configuration required by each specific library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi.
-On both datasets, as previously stated, only SparkSQL can complete the entire pipeline with the minimum configuration.
-Following closely is CuDF, which can handle the entire Patrol dataset with a laptop configuration, but needs a GPU, making it more expensive.
-Further, Vaex stands out on Patrol as one of the libraries that require the least resources, followed closely by SparkPD.
-ModinR registers a performance on Taxi similar to its Dask counterpart, but scales much better on Patrol.
-In fact, while Dask engine uses a centralized scheduler to share work across multiple cores, Ray uses distributed bottom-up scheduling.
-Indeed, Modin can run on top of Dask, but it was originally built to work with Ray, and such integration is still more mature.
-DataTable maintains a relatively consistent performance on both datasets.
-Instead, despite excelling in other experiments, scalability appears to be a significant weakness for Polars.
-Finally, Pandas continues to show severe limitations.
-TABLE V: Minimum machine configuration for running the entire pipeline on incremental dataset samples. Laptop, Workstation, Server. See Table IV for the detailed configurations.
+In particular, we take into account three different machine configurations, whose CPU and RAM are incrementally increased as described in Table 4 to reproduce typical configurations for laptops, workstations, and servers.
+Figure 6 shows the performance of the libraries for running the entire pipeline on incremental samples of Taxi.
+Table 5 outlines instead the minimum configuration required by each library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi.
+Since CuDF scales according to GPU memory, it is not included in this experiment.
+If the dataset fits in the GPU, CuDF is able to efficiently handle it and complete the pipeline.
+However, this dependency does not make it a good candidate for dealing with very large datasets on machines with limited GPU resources.
+Overall, SparkSQL clearly stands out as the best performer, being the only library capable of handling both datasets in their entirety on a laptop configuration.
+This capability is due to the use of disk spillover, which offloads data from RAM to disk storage, when the memory limit is reached during computations, ensuring continuous operation without exceeding memory capacity.
+On the other hand, SparkPD suffers instead from overhead due to internal conversions between Pandas and Spark dataframes and increased JVM memory usage, causing out-of-memory errors on the laptop configuration for both Patrol and Taxi.
+DataTable stands out on Patrol as one of the libraries that require the least resources, followed closely by ModinR and Pandas.
+Indeed, it is one of the libraries that consumes the least memory on Patrol; further, it does not require to split the dataframe into multiple pieces, which can lead to errors due to communication and information exchange.
+ModinR demonstrates better scalability across all machine configurations compared to its Dask counterpart.
+The Dask engine uses a centralized scheduler to distribute data across multiple cores, which results in higher memory consumption and leads to out-of-memory (OOM) issues more easily.
+In contrast, Ray can scale far beyond Dask due to its distributed task scheduling scheme, specifically employing a distributed bottom-up scheduling approach.
+This optimizes resource utilization and minimizes overhead by initiating tasks at the lowest level of computation and aggregating results upwards.
+Notably, Modin was originally built to work on top of Ray, making this integration more mature and optimized.
+Despite its excellent performance in other experiments, scalability appears to be a significant weakness for Polars.
+This limitation arises from its in-memory execution model, which requires all data to be loaded into RAM.
+While this approach offers speed and efficiency for smaller datasets, it becomes a bottleneck when handling larger datasets, limiting its scalability.
+Finally, Pandas confirms its well-known severe limitations about scalability, being the only library not able to complete the pipeline on Taxi even on a server configuration (Figure 6(c)).
+Table 5. Minimum machine configuration for running the entire pipeline on incremental dataset samples. Laptop, Workstation, Server. See Table 4 for the detailed configurations.
 Patrol Taxi
@@
 # Rows (M) 0.2 1.2 4.5 6.7 13.5 20.2 27 0.7 3.8 11.5 19.2 38.8 57.7 77
-Pandas LP LP WS SV SV LP WS SV
-SparkPD LP LP LP LP LP WS WS WS WS WS WS
+Pandas LP LP LP LP WS WS WS LP LP WS WS SV SV
+SparkPD WS WS WS WS WS WS WS WS WS WS WS WS WS WS
 SparkSQL LP LP LP LP LP LP LP LP LP LP LP LP LP LP
-ModinD LP LP WS WS WS SV SV LP LP WS WS SV
-ModinR LP LP LP LP WS WS SV LP LP WS WS SV
-Polars LP LP LP WS WS WS SV LP LP WS
-cuDF LP LP LP LP LP LP LP LP LP LP WS WS WS WS
-Vaex LP LP LP LP LP WS WS LP LP LP WS WS WS WS
-Datatable LP LP LP WS SV SV SV LP LP LP WS WS WS SV
-V Related Work
+ModinD LP LP LP WS SV SV SV LP LP LP WS SV
+ModinR LP LP LP LP WS WS WS LP LP LP LP WS WS WS
+Polars LP LP LP WS WS WS WS LP LP WS WS WS SV SV
+Vaex LP LP LP WS WS WS WS LP LP LP WS WS WS WS
+Datatable LP LP LP LP LP WS WS LP LP LP LP WS WS WS
+4.4. Performance on the TPC-H queries
+{mdframed}
+SummaryCuDF consistently achieves the best performance across all queries, while Polars significantly outperforms other CPU-only dataframe libraries.
+TPC-H [^27] https://www.tpc.org/tpch/default5.asp is a decision support benchmark consisting of a suite of business-oriented ad-hoc queries and concurrent data modifications (Pöss and Floyd, 2000).
+It is widely adopted to compare end-to-end database systems (Dreseler et al., 2020), allowing to comprehensively evaluate their performance and efficiency through the execution of complex queries on large volumes of data.
+Consistently with previous related work [^28] https://pola.rs/posts/benchmarks/ exploiting TPC-H queries to assess the performance of a subset of dataframe libraries (see Section 5 for more details), we also adopt this benchmark to provide further support to the outcomes of the comparative analysis presented above.
+In particular, we take as a reference a publicly available translation of the TPC-H queries into Pandas API operations [^29] https://github.com/Bodo-inc/Bodo-examples, replicating it for the other libraries, then use Bento for their seamless execution.
+Figure 7 illustrates the performance of each library on the 22 queries of the TPC-H 10GB benchmark.
+We select a scale factor of 10, because it represents the largest dataset that can be processed by the 40 GB RAM of our GPU, which therefore would not be able to handle a scale factor of 100.
+Overall, CuDF consistently achieves the best performance by leveraging GPU power.
+Among the CPU-only libraries, Polars is instead the clear winner.
+SparkSQL significantly outperforms Pandas on several queries, registering notable results especially on q06, where it is even faster than Polars (note that this query selects line items shipped within a year interval, reducing the dataset by 85%).
+Pandas is not the worst performer, but on most queries it does not achieve good results.
+Modin (which registers the same results with both engines, Dask and Ray) shows diverse performance instead, strongly dependent on the query.
+As mentioned above, SparkPD suffers from the additional latency due to the translation of Pandas operations into Spark execution plans, making it one of the worst performers.
+DataTable is one of the worst performers too, since it is slow in grouping operations (as also testified by their own benchmark [^30] https://h2oai.github.io/db-benchmark/) and its API only supports joins on columns with unique values, requiring therefore to switch to the default Pandas API (e.g., q09).
+Finally, Vaex is by far the worst performer overall, as it is significantly slow in grouping operations (as also pointed out in the previous sections) and lacks support for multi-column joins.
+5. Related Work
 To the best of our knowledge, our paper presents the first rigorous and extensive experimental comparison of existing dataframe libraries on data preparation tasks.
-In fact, public wisdom about such libraries is mostly scattered across several not peer-reviewed sources [16, 17, 18, 19, 20], which generally compare the performance of few libraries (e.g., Pandas vs PySpark vs Polars) on a handful of operations.
+In fact, public wisdom about such libraries is mostly scattered across several not peer-reviewed sources (Schmitt, 2020; Alexander, 2023; Pinner, 2023), which generally compare the performance of few libraries (e.g., Pandas vs PySpark vs Polars) on a handful of operations.
 The major claims by such analysis are mostly confirmed by our evaluation; nevertheless, despite often providing some useful insights, they only offer a partial and very fragmented knowledge, far from a complete and detailed overview covering all libraries and operations.
-In literature, papers presenting new dataframe libraries or dataframe-based tools usually include a comparison against some dataframe libraries.
-It is the case for the AFrame data analysis package [56] (later extended to PolyFrame [55]), designed to efficiently manage large-scale dataframes using SQL++ queries.
-AFrame is evaluated against Pandas, Spark, and Pandas on Ray (i.e., ModinR) through a micro-benchmark using the synthetic Wisconsin benchmark dataset [57] to assess acceleration and scalability across distributed environments.
+In literature, when a paper introduces a new dataframe library or dataframe-based tool, it usually includes a comparison against some dataframe libraries.
+It is the case for the AFrame data analysis package (Sinthong and Carey, 2019) (later extended to PolyFrame (Sinthong and Carey, 2021)), designed to efficiently manage large-scale dataframes using SQL++ queries.
+AFrame is evaluated against Pandas, Spark, and Pandas on Ray (i.e., ModinR) through a micro-benchmark using the synthetic Wisconsin benchmark dataset (DeWitt, 1993) to assess acceleration and scalability across distributed environments.
 Similarly, Petersohn et al.
-[21] evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas.
-However, the mentioned comparisons share significant limitations, since data preparation is not the main topic of the analysis and only a relatively small set of operations is considered.
+(Petersohn et al., 2021) evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas.
+However, these comparisons share significant limitations, since data preparation is not their main focus and only a relatively small set of libraries and operations is considered.
 Further, dataframe libraries are sometimes compared in data science benchmarks.
-For instance, TPC-H [58], a popular decision support benchmark including a suite of business-oriented ad hoc queries and concurrent data modifications, is used by Polars [24] to compare against Pandas, PySpark, Dask, Vaex, and DuckDB [59].
-Similarly, H2O (creator of DataTable[60] uses a set of queries to asses the scalability of group by and join operations on popular database and dataframe-like systems.
-Nevertheless, such relational benchmarking systems notoriously offer very limited insights about real-world workflows executed on dataframes [22].
-Sanzu [61] is instead designed to evaluate five popular data science frameworks and systems (R, Anaconda [62], Dask, PostgreSQL [63], and PySpark) on data processing and analysis tasks.
+For instance, as previously mentioned in Section 4.4, it is the case for TPC-H (Pöss and Floyd, 2000), which is used by Polars  [^31] https://pola.rs/posts/benchmarks/ to compare against Pandas, PySpark, Dask, Modin, and DuckDB (Raasveldt and Mühleisen, 2019).
+While the outcome of their analysis is consistent with our findings, their evaluation (still performed by selecting a scale factor of 10) only covers the first 7 queries out of the entire batch of 22 covered by our evaluation.
+Similarly, H2O (creator of DataTable) uses a set of queries to asses the scalability of popular database and dataframe-like systems [^32] https://h2oai.github.io/db-benchmark only limited to group by and join operations.
+Indeed, due to their different focus, such relational benchmarking systems notoriously offer very limited insights into real-world data preparation workflows executed on dataframes (Rehman and Elmore, 2022).
+Note that for producing a consistent a analysis we do not take into account tools that are not based on the dataframe data structure, as the aforementioned DuckDB or similar SQL-based solutions.
+For instance, Grizzly (Kläbe and Hagedorn, 2021) aims to address Pandas limitations in scalability by translating dataframes to SQL code, while PyFroid translates from the Pandas API to the DuckDB API to efficiently scale Pandas workloads on a commodity workstation, using top-voted Kaggle notebooks for the evaluation.
+Sanzu (Watson et al., 2017) is designed to evaluate five popular data science frameworks and systems (R, Anaconda [^33] https://www.anaconda.com, Dask, PostgreSQL [^34] https://www.postgresql.org, and PySpark) on data processing and analysis tasks.
 Sanzu is composed of a micro-benchmark, to test basic operations in isolation on synthetic datasets of different sizes, and a macro-benchmark, to evaluate series of operations representing concrete data science scenarios on real-world datasets.
-FuzzyData [22] is a workflow generation system that allows to compare dataframe-based APIs on worflows composed of a small subset of operations.
+FuzzyData (Rehman and Elmore, 2022) is a workflow generation system that allows to compare dataframe-based APIs on workflows composed of a small subset of operations.
 Finally, Shanbhag et al.
-[23] measure the energy consumption of 21 dataframe processing tasks on Pandas, Vaex, and Dask.
-VI Key Takeaways
-Our evaluation confirmed the several known limitations of Pandas.
-However, despite the promises of the many libraries designed to overtake it, there is no silver bullet for data preparation of tabular data with dataframes on a single machine.
+(Shanbhag and Chimalakonda, 2023) measure the energy consumption of 21 dataframe processing tasks on Pandas, Vaex, and Dask.
+(a) Queries 1 to 11.
+(b) Queries 12 to 22.
+Figure 7. Performance of the dataframe libraries on the TPC-H 10GB queries.
+6. Key Takeaways
+Our evaluation confirmed the well-known limitations of Pandas.
+However, despite the abundance of alternative dataframe libraries designed to overtake it, there is no silver bullet to perform data preparation of tabular data with dataframes on a single machine.
+The right choice usually depends on many factors, such as the size and the features of the dataset at hand, the configuration of the underlying machine, and the kind of operations to carry out.
 We can help data practitioners to decide which library to use by answering the following questions.
-VI-A Does the dataset fit in the GPU memory?
+Does the dataset fit in the GPU memory?
 If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API.
-However, CuDF lacks of an optimizer, thus libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations [20].
-VI-B No complete pipeline execution?
-If a user is only interested in an individual stage (I/O, EDA, data transformation, or data cleaning), DataTable stands out as the fastest library for I/O operations—for other tasks it might not be that mature, with many preparators to implement manually.
-Polars emerges as the best performer for EDA.
-However, for Data Transformation and Data Cleaning, SparkSQL and Vaex show the best results.
-VI-C What is the size of the datasets?
-With small datasets ($<500k$ rows), all of the libraries show similar average runtimes.
-Even if Polars is often recommended for such a scenario, Pandas (especially Pandas2) appears to be the most reasonable choice instead, since their performance does not differ substantially, at least not enough to learn a new API and replace entire parts of the pipeline [18].
-In general, if memory and CPU usage are limited, Polars has the best overall performance.
-For datasets of medium size (from 2 to 19 million rows) with dense numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed.
+However, CuDF lacks of an optimizer; thus, libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations [^35] https://pola-rs.github.io/polars/user-guide/misc/alternatives/.
+Individual stage or complete pipeline?
+If a user is only interested in an individual stage, i.e., input/output (I/O), explorative data analysis (EDA), data transformation (DT), or data cleaning (DC), DataTable stands out as the fastest library for I/O operations, while for other tasks it might not be that mature.
+Further, it shows the least compatibility with Pandas, requiring to implement many preparators manually.
+Polars emerges as the best performer for EDA, while SparkSQL represents a consistent solution for DT and DC, also recording the best performance among all CPU-only libraries when running the entire pipeline.
+What is the size of the datasets?
+For small datasets (less than 500k rows), all of the libraries show similar average runtimes.
+Even if Polars is often recommended for such a scenario, and indeed registers the best overall performance if memory resources are not limited, Pandas might still be the most reasonable choice, since their performance does not differ substantially—at least, not enough to learn a new API and replace entire parts of the pipeline (Pinner, 2023).
+In the case of datasets of medium size (from 2 to 19 million rows) with dense numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed.
 On the other hand, for sparse datasets with many null values (e.g., Loan), Polars and SparkPD show good performance.
 Finally, for datasets with many string columns (e.g., 25% Patrol), SparkSQL appears to be the best choice.
-For larger datasets (over 20 million rows), SparkSQL stands out as the best option [64].
+For larger datasets (over 20 million rows), SparkSQL stands out as the best option (Tang, 2023).
 Furthermore, when the pipeline requires dealing with heavy-duty column operations, particularly in datasets that have a multitude of string columns and a good amount of null values (e.g., Patrol), Vaex seems to be the best move—and there is no need to set up any environment, as required for PySpark.
-VII Conclusion
-We presented an experimental comparison of the most used dataframe-based libraries to help practitioners in selecting the best suitable to run their data preparation tasks on a single machine.
-To guarantee reasonably comparable results and increase usability, we developed Bento, a general framework for assessing library performance on four major data preparation stages: Input/Output (I/O), Exploratory Data Analysis (EDA), Data Transformation (DT), and Data Cleaning (DC).
-Our contribution includes a thorough analysis of performance of the most used dataframe libraries on heterogeneous datasets, that ranges in size, complexity, and characteristics and the relative most voted pipeline from Kaggle.com.
-The results of our experiments show that there is no clear winner among the libraries, thus we summarized useful takeaways for supporting the selection of the right library for a task and dataset at hand.
-As future work, we are planning to compare the libraries in a distributed setting and to study how machine learning can support users in suggesting optimal library combinations based on datasets, preparators, and previous executions.
-Acknowledgment
-We extend our appreciation to the master’s students of the 2021/2022 academic year in the Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions to testing the libraries.
-This work was partially supported by the Department of Engineering “Enzo Ferrari” within the FARD-2022 (FAR2022DIP_DIEF - Prot.
-n.
-2985 del 27/07/2022) and FARD-2023 (FAR_DIP_2023_DIEF-SIMONINI) projects.
+7. Conclusion
+We presented a comprehensive experimental comparison of the most used dataframe libraries to support practitioners in the selection of the most suitable solution to carry out their data preparation tasks on a single machine.
+To guarantee comparable results and increase usability, we developed Bento, a general framework for assessing the performance of dataframe libraries on four major data preparation stages: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC).
+We exploited Bento to perform a thorough comparative analysis of the libraries on heterogeneous datasets (varying in size, complexity, and features) previously adopted in literature and publicly available on Kaggle.
+In particular, we relied on the most voted Kaggle notebooks to assess the performance of the libraries on real-world data preparations pipelines validated by a large community.
+Further, we also compare their performance on the popular TPC-H benchmark to support our findings.
+The results of our experiments show that there is no clear winner among the libraries, whose performance is strongly dependent on the size and features of the dataset at hand, on the specifications of the underlying machine, and on the kind of operations to carry out.
+Thus, we provided the readers with useful takeaways to support them in the selection of the right library for the data preparation task at hand.
+As future work, we plan to compare the libraries in a distributed environment and study the possible adoption of machine learning to suggest optimal library combinations based on datasets, tasks, hardware, and previous executions.
+Acknowledgements
+We extend our appreciation to the master’s students of the 2021/2022 Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions in testing the libraries.
+We warmly thank Leonardo S.p.A., co-funding the PhD program of Angelo Mozzillo.
+Further, this work was partially supported by the “Enzo Ferrari” Engineering Department (University of Modena and Reggio Emilia) within the projects FARD-2022 and FARD-2023, and by MUR within the project “Discount Quality for Responsible Data Science: Human-in-the-Loop for Quality Data” (code 202248FWFS).

### v3

@@
-and Reggio Emilia, Italy
+Evaluation of Dataframe Libraries for Data Preparation on a Single Machine
 Abstract.
@@
 In this paper, we evaluate the most popular Python dataframe libraries in general data preparation use cases to assess how they perform on a single machine.
-To do so, we employ 4 real-world datasets and pipelines with heterogeneous features, covering a variety of scenarios.
+To do so, we employ 4 real-world datasets with heterogeneous features, covering a variety of scenarios, and the TPC-H benchmark.
 The insights gained with this experimentation are useful to data scientists who need to choose which of the dataframe libraries best suits their data preparation task at hand.
-In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when data fits in RAM and there is no need for complete compatibility with Pandas API, Polars is the go-to choice thanks to its in-memory execution and query optimizations; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multithread execution and a query optimizer) proves to be the best options.
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 28th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06
+In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when data fits in RAM and there is no need for complete compatibility with Pandas API, Polars is the go-to choice thanks to its in-memory execution and query optimizations; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multithread execution and a query optimizer) proves to be the best option.
+copyright: rightsretaineddoi: https://doi.org/10.48786/edbt.2025.27isbn: 978-3-89318-098-1conference: 28th International Conference on Extending Database Technology (EDBT); 25th March-28th March, 2025; Barcelona, Spainjournalyear: 2025
 1. Introduction
@@
 Thus, guaranteeing high data quality is fundamental to ensure the reliability of analysis and avoid undesired additional costs (Haug et al., 2011).
-Oftentimes data is heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as missing or duplicate values (Fan, 2015).
+Data is often heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as missing or duplicate values (Fan, 2015).
 Due to that, data scientists dedicate considerable time and resources to perform data preparation (Fernandes et al., 2023).
@@
 Libraries for data preparation typically offer a flexible functional interface that allows to conveniently compose pipelines.
-In particular, Pandas (McKinney, 2010) is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators (Petersohn et al., 2020; Jindal et al., 2021; Psallidas et al., 2022; Lee et al., [n.
-d.]).
+Pandas (McKinney, 2010) is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators (Petersohn et al., 2020; Jindal et al., 2021; Psallidas et al., 2022; Lee et al., 2021).
 This library has become so popular in the data science ecosystem that even the popularity of Python itself has been sometimes attributed to its wide adoption (Claburn, 2017).
-Despite its success, Pandas notoriously presents some severe limits (Emani et al., 2024; Kläbe and Hagedorn, 2021).
+Despite its success, Pandas notoriously presents multiple severe limits (Emani et al., 2024; Kläbe and Hagedorn, 2021).
 Firstly, it has not been designed to work with large datasets efficiently.
@@
 However, these libraries come with heterogeneous features (e.g., lazy evaluation, GPU support, etc.), making it hard for data scientists to navigate among them and choose the most suitable solution.
-A major reason for this confusion is the lack of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario.
+In particular, data scientists lack the support of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario.
 Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons (Schmitt, 2020; Alexander, 2023; Pinner, 2023) (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature (Petersohn et al., 2021; Rehman and Elmore, 2022; Shanbhag and Chimalakonda, 2023) either focus on orthogonal aspects or only cover small subsets of libraries and preparators.
@@
 Hence, the faster the employed library is, the less idle time they have.
-To support data scientists in the choice of the dataframe library that can speedup their workflow, we present the first extensive experimental comparison of Pandas and its most popular dataframe-based alternatives, namely PySpark, Modin (with both Dask and Ray engine), Polars, CuDF, Vaex, and DataTable.
+To support data scientists in the choice of the library that can speedup their workflow, we present the first extensive experimental comparison of Pandas and its most popular dataframe-based alternatives, namely PySpark, Modin, Polars, CuDF, Vaex, and DataTable.
 For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library.
 Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries and enables the execution in a containerized environment using Docker.
-This process is streamlined with the help of a convenient configuration file.
-To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, previously adopted in related literature on data preparation (Hameed and Naumann, 2020) and dataframe libraries (Petersohn et al., 2021).
-All these datasets are collected from Kaggle [^2] https://www.kaggle.com, the most popular platform for data science challenges.
+This process is streamlined with the help of a convenient configuration file [^2] https://github.com/dbmodena/bento/tree/master#write-a-test-file.
+To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, previously adopted in related work on data preparation (Hameed and Naumann, 2020) and dataframe libraries (Petersohn et al., 2021).
+All these datasets are collected from Kaggle (Kaggle, 2024), the most popular platform for data science challenges.
 This allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and have been validated by the world’s largest data science community.
-We identify four essential data preparation stages in the pipelines: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC).
+Moreover, we further validate our findings by testing the different libraries with the TPC-H benchmark (Pöss and Floyd, 2000).
+We identify four essential data preparation stages: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC).
 Thus, we evaluate the performance of the libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
@@
 In fact, compared to specific data preparation tools (Hameed and Naumann, 2020), dataframes offer greater simplicity and customization, enabling users to manipulate the data by writing Python code without the effort of learning to use a new tool.
-Pandas  [^4] https://github.com/pandas-dev/pandas has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language (McKinney, 2010)) and provides efficient and expressive data structures optimized to work with structured datasets.
+Pandas (Pandas, 2024a) has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language (McKinney, 2010)) and provides efficient and expressive data structures optimized to work with structured datasets.
 Pandas has established itself over time as the standard library for data manipulation and analysis.
@@
 To address such limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing.
-The selection of the Python libraries to include in our evaluation was based on three main criteria, requiring them to: (i) be based on the dataframe data structure (excluding therefore SQL-based solutions such as DuckDB (Raasveldt and Mühleisen, 2019)); (ii) be compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the de facto standard, already aim at Pandas-compatibility for their API [^5] https://modin.readthedocs.io/en/latest or describe the parallelism between their API and the Pandas API [^6] https://datatable.readthedocs.io/en/latest/manual/comparison_with_pandas.html); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support.
+The selection of the Python libraries to include in our evaluation was based on three main criteria, requiring them to: (i) be based on the dataframe data structure (excluding therefore SQL-based solutions such as DuckDB (Raasveldt and Mühleisen, 2019)); (ii) be compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the de facto standard, already aim at Pandas-compatibility for their API (Modin, 2024a) or describe the parallelism between their API and the Pandas API (DataTable, 2024a)); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support.
 The selected libraries are: Pandas, PySpark, Modin, Polars, CuDF, Vaex, and DataTable.
@@
 Deploy on cluster
-[…]
+Native language Python Scala Python Rust C/C++ C/Python C++/Python
 Licence 3-Clause BSD Apache 2.0 Apache 2.0 MIT Apache 2.0 MIT Mozilla Public 2.0
-[…]
+Other requirements SparkContext Ray/Dask CUDA
 Considered version 2.2.1 3.5.1 0.29.0 0.20.23 24.04.01 4.17.0 1.1.0
@@
 For each dataframe library, Table 1 lists which of these features it implements, along with information about its implementation and the version considered here.
-PySpark [^7] https://github.com/apache/spark is the Python API for Apache Spark, which is primarily implemented in Scala.
-Spark allows to work with distributed collections of Java or Python objects partitioned across a cluster, known as Relational Distributed Datasets (RDDs) (Armbrust et al., 2015).
-PySpark dataframes are RDDs consisting of Row objects: they can be seen as relational tables, on which relational operators can be applied.
-Further, they support lazy evaluation, relying on the Catalyst optimizer (Armbrust et al., 2015).
+PySpark (Apache Spark, 2024a) is the Python API for Apache Spark, which is primarily implemented in Scala.
+PySpark DataFrame is a type of Dataset (i.e., a distributed collection of data) organized into named columns, resembling a relational table and supporting relational operators (Apache Spark, 2024d).
+PySpark also supports lazy evaluation, relying on the Catalyst optimizer (Armbrust et al., 2015) and a disk spillover mechanism that automatically offloads data from RAM to disk when memory limits are reached (Apache Spark, 2024e).
+This allows PySpark to process datasets that exceed the machine’s physical memory capacity.
 While Spark owes its popularity to the capability of processing large-scale datasets on a cluster of machines, standalone (i.e., single-machine) mode is not only supported, but also surprisingly fast in multiple scenarios, as shown in Section 4.
-PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs for working with structured data; (ii) Pandas on Spark (denoted as SparkPD and previously known as Koalas [^8] https://github.com/databricks/koalas), which enables (by adding an index to the conventional Spark dataframe) to distribute Pandas workloads across multiple nodes without requiring modifications to the original Pandas code for most API functions ($\sim$80%).
-Table 2.
-Compatibility of dataframe libraries with Pandas API.
-(✓✓) fully matches Pandas interface; (✓) different interface; ($\circ$) missing from the API, but implemented by us to the best of our efforts.
-Preparator Pandas API SparkPD SparkSQL Modin Polars CuDF Vaex DataTable
-I/O load dataframe read_csv ✓✓ ✓✓ ✓✓ ✓✓
-output dataframe to_csv ✓✓ ✓✓ ✓✓ ✓✓
-EDA locate missing values (nulls) isna ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
-locate outliers percentile ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
-search by pattern str.contains ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
-sort values sort ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
-get columns list columns ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
-get columns types dtypes ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
-get dataframe statistics describe ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
-query columns query ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
-DT cast columns types astype ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
-delete columns drop ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
-rename columns rename ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
-pivot pivot_table ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
-calculate column using expressions apply columnwise ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
-join dataframes merge ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ $\circ$
-one hot encoding get_dummies ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ $\circ$
-categorical encoding cat.codes ✓✓ ✓✓ ✓✓ $\circ$
-group dataframe groupby ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
-DC change date & time format to_datetime ✓✓ ✓✓ $\circ$ ✓✓ $\circ$ $\circ$
-delete empty and invalid rows dropna ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
-set content case str.lower ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
-normalize numeric values round ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
-deduplicate rows drop_duplicates ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
-fill empty cells fillna ✓✓ ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
-replace values occurrences replace ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
-edit & replace cell data apply rowise ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓
-Modin [^9] https://github.com/modin-project/modin is a Python library that provides a parallel alternative to Pandas (Petersohn et al., 2020, 2021).
+PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs to work with structured data; (ii) Pandas on Spark (denoted as SparkPD and previously known as Koalas (Databricks, 2024)), which enables (by adding an index to the conventional Spark dataframe) to distribute Pandas workloads across multiple nodes without requiring modifications to the original Pandas code for most API functions ($\sim$80%).
+Modin (Modin, 2024b) is a Python library that provides a parallel alternative to Pandas (Petersohn et al., 2020, 2021).
 Modin adopts the Pandas data format as the default storage layer and employs a set of 15 core operators to simplify Pandas functions and build its own Pandas-like API.
@@
 Each partition is then processed independently by the execution engine: Dask or Ray (we consider both solutions, denoted as ModinD and ModinR, respectively).
-Dask  [^10] https://github.com/dask/dask is a Python library for distributed computing, which allows working with large distributed Pandas dataframes (Rocklin, 2015).
+Dask (Dask, 2024b) is a Python library for distributed computing, which allows working with large distributed Pandas dataframes (Rocklin, 2015).
 It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory, and its scheduler provides flexibility to the execution.
-While we also considered the inclusion of Dask as an independent library, we do not report its results since we found it to perform better in combination with Modin (as Dask is not well suitable for a single machine) and it only covers $\sim$55of the Pandas API (while Modin $\sim$90%).
-Ray  [^11] https://github.com/ray-project/ray is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Apache Arrow [^12] https://github.com/apache/arrow (a language-independent columnar memory format, recently adopted even by Pandas  [^13] https://pandas.pydata.org/docs/dev/whatsnew/v2.0.0.html) as data format (Moritz et al., 2018).
+While we also considered the inclusion of Dask as an independent library, we do not report its results since we found it to perform better in combination with Modin (as Dask is not well suitable for a single machine) and it only covers $\sim$55of the Pandas API (while Modin covers $\sim$90%) (Modin, 2024c).
+Ray (Ray, 2024b) is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Apache Arrow (Apache Arrow, 2024a) (a language-independent columnar memory format, recently adopted even by Pandas (Pandas, 2023)) as data format (Moritz et al., 2018).
 Unlike Dask, Ray does not have built-in primitives for partitioned data.
 While the two engines serve different use cases, their primary goal is similar: optimizing resource usage by changing how data is stored and Python code is executed.
-Polars [^14] https://github.com/pola-rs/polars is a Python library written in Rust and built on top of arrow2 (the Rust implementation of Arrow format).
-The adoption of Arrow as underlying data structure confers efficient data processing capabilities to Polars, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
+Polars (Polars, 2024d) is a Python library written in Rust and built on top of arrow2, the Rust implementation of the Arrow format.
+The adoption of Arrow as the underlying data structure provides Polars with efficient data processing capabilities, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
 Polars does not use an index, but each row is indexed by its integer position in the table.
@@
 Polars optimizes queries through early filters and projection pushdown.
-Moreover, it supports both eager and lazy execution.
+Moreover, it supports both eager and lazy evaluation.
 The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can run in batches, lightening the load on memory and CPU, hence allowing to process bigger datasets (even larger than memory) in a faster manner.
-CuDF [^15] https://github.com/rapidsai/cudf is a component of the NVIDIA RAPIDS [^16] https://rapids.ai framework.
+CuDF (NVIDIA, 2024a) is a component of the NVIDIA RAPIDS framework.
 Written in C/C++ and CUDA (hence only compatible with NVIDIA GPUs), it offers a Pandas-like API to run general-purpose data science pipelines on GPUs, leveraging their computational power for accelerating data processing.
@@
 Thus, it can perform efficient and high-performance computations, although it does not provide any optimization strategy for the execution of the pipelines.
-Vaex [^17] https://github.com/vaexio/vaex is a Python library [^18] Since the last commit on GitHub is dated about 10 months ago, the project might not be actively maintained at the moment.
+Vaex (Vaex, 2024c) is a Python library [^4] Since the last commit on GitHub is dated about one year ago, the project might not be actively maintained at the moment.
 designed to handle extremely large tabular datasets (such as astronomical catalogs), producing fast statistical analysis and visualization.
-Supported by its vaex-core extension, written in C, it achieves memory optimization through streaming algorithms, memory-mapped files, and a zero-copy policy, allowing to explore datasets exceeding the available memory.
+Supported by its vaex-core extension, written in C, Vaex achieves memory optimization through streaming algorithms, memory-mapped files, and a zero-copy policy, supporting the exploration of datasets that exceed the available memory.
 Vaex enables efficient column-wise operations by wrapping a series of NumPy arrays as columns.
@@
 Vaex can process random data subsets and export data in random order for efficient resource utilization.
-Further, it also supports multithreading for faster computation on multi-core CPUs.
-DataTable [^19] https://github.com/h2oai/datatable is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns).
+Moreover, it also supports multithreading for faster computation on multi-core CPUs.
+DataTable (H2O, 2024) is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns).
 DataTable, which uses a native-C implementation for all data types, is developed to support column-oriented data storage, optimizing data access and processing.
@@
 Datasets
-For a comprehensive evaluation of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets covering a variety of sizes, complexities, and features: (i) Athlete  [^20] https://www.kaggle.com/datasets/heesoo37/120-years-of-olympic-history-athletes-and-results, collecting 0.2M records about the results achieved by the athletes through 120 years of Olympics; (ii) Loan  [^21] https://www.kaggle.com/datasets/wordsforthewise/lending-club, containing 2M rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol  [^22] https://www.kaggle.com/datasets/faressayah/stanford-open-policing-project, composed of 27M records about 11 years of traffic stops by California law enforcement agencies (Pierson et al., 2020); (iv) Taxi  [^23] https://www.kaggle.com/competitions/nyc-taxi-trip-duration, containing 77M records about taxi trips in New York City in 2015.
+Table 2. Features of the selected datasets.
+Athlete Loan Patrol Taxi CSV Size (GB) 0.03 1.6 6.7 10.9 # Rows (×$\times 10^{6}$) 0.2 2 27 77 # Columns 15 151 34 18 # Num - Str - Bool 5-10-0 113-38-0 5-27-2 15-3-0 % Null 9% 31% 22% 0% Str Len Range (1, 108) (1, 3988) (1, 2293) (1, 19)
+For a comprehensive evaluation of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets covering a variety of sizes, complexities, and features: (i) Athlete (Kaggle, 2018), collecting 0.2M records about the results achieved by the athletes through 120 years of Olympics; (ii) Loan (Kaggle, 2019a), containing 2M rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol (Kaggle, 2019b), composed of 27M records about 11 years of traffic stops by California law enforcement agencies (Pierson et al., 2020); (iv) Taxi (Kaggle, 2017), containing 77M records about taxi trips in New York City in 2015.
 A summary of their main features is displayed in Table 2.
+Note that we compute the percentage of missing values as the number of null cells over all cells in the dataframe.
 Collected from Kaggle, all of these datasets had previously been adopted in related work about data preparation and dataframe libraries.
@@
 Instead, Athlete was exploited to analyze how different preparators are supported by commercial data preparation tools (Hameed and Naumann, 2020).
-Table 3. Features of the selected datasets.
-Athlete Loan Patrol Taxi CSV Size (GB) 0.03 1.6 6.7 10.9 # Rows (×$\times 10^{6}$) 0.2 2 27 77 # Columns 15 151 34 18 # Num - Str - Bool 5-10-0 113-38-0 5-27-2 15-3-0 % Null 9% 31% 22% 0% Str Len Range (1, 108) (1, 3988) (1, 2293) (1, 19)
 Data Preparation Pipelines
-For each dataset, we selected from Kaggle the most voted notebook that proposed a non-trivial solution for the task at hand (i.e., Kaggle competitions) and extracted its data preparation pipeline—typically the first part of the notebook, preceding the definition of a machine learning model.
+Table 3.
+Compatibility of dataframe libraries with Pandas API.
+(✓✓) fully matches Pandas interface; (✓) different interface; ($\circ$) missing from the API, but implemented by us to the best of our efforts.
+Preparator SparkPD SparkSQL Modin Polars CuDF Vaex DataTable
+I/O load dataframe (read) ✓✓ ✓✓ ✓✓ ✓✓
+output dataframe (write) ✓✓ ✓✓ ✓✓ ✓✓
+EDA locate missing values (isna) ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
+locate outliers (outlier) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
+search by pattern (srchptn) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
+sort values (sort) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
+get columns list (getcols) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
+get columns types (dtypes) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
+get dataframe statistics (stats) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
+query columns (query) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
+DT cast columns types (cast) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
+delete columns (drop) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
+rename columns (rename) ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
+pivot table (pivot) ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
+calculate column using expressions (calccol) ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
+join dataframes (join) ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ $\circ$
+one hot encoding (onehot) ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ $\circ$
+categorical encoding (catenc) ✓✓ ✓✓ ✓✓ $\circ$
+group dataframe (group) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
+DC change date & time format (chdate) ✓✓ ✓✓ $\circ$ ✓✓ $\circ$ $\circ$
+delete empty and invalid rows (dropna) ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
+set content case (setcase) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
+normalize numeric values (norm) ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
+deduplicate rows (dedup) ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
+fill empty cells (fillna) ✓✓ ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
+replace values occurrences (replace) ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
+edit & replace cell data (edit) ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓
+For each dataset, we selected three notebooks among the top-voted Kaggle entries that proposed a non-trivial solution for the task at hand (i.e., Kaggle competitions) and extracted its data preparation pipeline—typically the first part of the notebook, which precedes the definition of a machine learning model.
 Overall, the pipelines employ a set of 27 preparators (reported in Table 3) that can be clustered into four main stages:
   • Input/output (I/O): preparators to handle the input/output of data in various formats, such as databases, CSV or Parquet files, Web APIs, etc.
-  • Exploratory data analysis (EDA): preparators for supporting the exploratory analysis of data features, to better understand the data at hand and detect errors or anomalies.
+  • Exploratory data analysis (EDA): preparators to support the exploratory analysis of data features, to better understand the data at hand and detect errors or anomalies.
   • Data transformation (DT): preparators for transforming the data to make it more suitable for analysis, such as categorical encoding, join, aggregation, etc.
@@
 […]
-Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each stage (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline.
-Finally, our framework also supports the pipeline execution in Docker containers, which allow to specify core and memory usage and isolate Python environments to prevent package conflicts.
+Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each of the four stages (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline.
+Finally, our framework also supports the execution of the pipeline in Docker containers, which allow to specify core and memory usage and isolate Python environments to prevent package conflicts.
 Hardware and Software
@@
 Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks.
-Thus, our comparison is designed to assess the performance of dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
+Therefore, our comparison is designed to assess the performance of dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
 (a) Athlete, EDA
@@
 (l) Taxi, DC
-Figure 1. Average runtime for each stage of the data preparation pipeline (EDA, DT, DC) on different datasets.
+Figure 1. Average speedup over Pandas computed for each stage (EDA, DT, DC) on the three data preparation pipelines.
 In particular, we organize our evaluation in multiple subsections, each designed to answer one of the following research questions:
@@
 {mdframed}
-Summary—For EDA, Polars is generally the best performer, up to 100 times faster than Pandas.
+Summary—For EDA, Polars is generally the best performer.
 For DT, if a GPU is available, CuDF generally outperforms other libraries.
-For DC, ModinR achieves notable results on the largest datasets.
-Finally, despite not supporting the Parquet format, DataTable is the fastest library for I/O.
-For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 shows the runtime for the execution of three main stages of the data preparation pipeline: EDA, DT, and DC.
+For DC, Vaex achieves notable results on the largest datasets.
+Finally, CuDF and Polars appear to be the best choices to read and write CSV files, respectively.
+For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 and Figure 2 show the average speedup over Pandas achieved by its alternatives in the execution of the three data preparation pipelines per dataset, focusing on the stages of EDA, DT, and DC.
+In particular, Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at the stage level when supported, Figure 2 shows the performance separately for each distinct preparator (i.e., we force the execution for each of them).
 The I/O stage is considered separately in Figures 4 and 4.
-Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at stage level when supported.
-Additionally, in Figure 2 we present the independent performance of each distinct preparator (i.e., we force the execution separately for each preparator), normalized with respect to Pandas (the de facto standard library for managing dataframes).
-For each preparator, the red line represents therefore Pandas as the baseline, while distinct markers denote the speedup for each library, which is defined as follows:
-$speedup=\frac{Time{\langle\textrm{Pandas},\textit{{prep}}\rangle}}{Time{\langle\textit{{lib}},\textit{{prep}}\rangle}}$
-where $Time{\langle\textit{{lib}},\textit{{prep}}\rangle}$ is the time required by the library lib to run the preparator prep.
-Thus, a value above the red line denotes that the library outperforms Pandas; otherwise, it takes more time than Pandas to run the preparator.
-The logarithmic scale enables to easily assess the order of magnitude of the improvement (if any) of each library over Pandas.
+The bars in Figure 1 and the markers in Figure 2 denote the speedup achieved over Pandas (the red line), defined as follows:
+$\textit{{speedup}}=\frac{Time{\langle\textit{{Pandas}},\textit{{prep/stage}}\rangle}}{Time{\langle\textit{{lib}},\textit{{prep/stage}}}\rangle}$
+[…]
+$\textit{{Impact}}=\frac{Time{\langle\textit{{dataset}},\textit{{prep}}\rangle}}{Time{\langle\textit{{dataset}},\textit{{stage}}\rangle}}\times 100$
+where $Time{\langle\textit{{dataset}},\textit{{prep}}\rangle}$ is the average time for each stage preparator and $Time{\langle\textit{{dataset}},\textit{{stage}}\rangle}$ is the sum of the average times for all preparators of the stage on the three dataset pipelines [^5] To minimize bias from particularly fast or slow libraries, both in the numerator and in the denominator time values below the 20th percentile and above the 80th percentile were excluded..
 In the next subsections, we analyze in detail the results obtained for each of the main data preparation stages, considering both the single preparators and the entire stages.
-Note that not all preparators have the same impact on the pipeline, but their weight depends on the time required for their execution (as we point out when analyzing the related results).
 (a) Athlete
@@
 (d) Taxi
-Figure 2. Speedup over Pandas for each preparator on different datasets.
-Explorative Data Analysis (EDA)
-For EDA, Polars clearly stands out as the best performer on all datasets but Patrol, where CuDF outperforms other dataframe libraries (Figures 1(a)-1(d)).
-In this stage, four preparators alone account for 95% of the total execution time: locate missing values (isna), locate outliers (outlier), search by pattern (srchptn), and sort values (sort).
-Note that Polars generally registers the best performance for two of them, locate missing values and locate outliers (Figure 2), where it is up to ${10^{4}}$ times faster than Pandas and 0.6 times faster than the second best performer, DataTable.
-In fact, Polars can quickly locate missing values by exploiting the specific metadata stored by Arrow arrays [^24] https://pola-rs.github.io/polars/user-guide/expressions/null/.
-Nevertheless, DataTable is the fastest in locating missing values on Patrol (Figure 2(c)), with a speedup of ${10^{4}}$ over Pandas and ${10^{3}}$ over Polars, thanks to its implementation of binary-search-based filtering.
-As highlighted by Table 3, Patrol is mostly composed of string attributes.
-For this type of data, Polars uses Arrow large_utf8 buffers, which are more expensive to filter than Python strings/chars, requiring therefore extra work.
-For the other two most relevant preparators, in Figure 2(c) we can observe that Vaex excels in searching by pattern, being about ${10^{3}}$ times faster than Pandas and 10 times faster than the second best performer, CuDF.
-Vaex tends to be very efficient in column-wise operations due to its internal columnar organization of the data.
-On the other hand, it is much less efficient when applying a row-wise preparator, for instance in locating outliers, where it is even about 100 times slower than Pandas.
-Conversely, when locating missing values, Vaex shows relatively good performance on dense datasets such as Taxi and Patrol, though it is among the less efficient on sparse datasets like Loan, where its performance is does not provide improvements over Pandas (Figure 2(b)).
-For sorting values, CuDF and Polars are the fastest libraries overall.
-SparkSQL shows remarkable advantages over Pandas as the datasets grow larger, while SparkPD maintains similar performance to the baseline.
-Note that in Athlete, Modin converts by default the dataframe back and forth to Pandas for the sorting operation on multiple columns, hence in this case it is the worst performer, up to 100 times slower than Pandas.
-Finally, for the remaining preparators, get dataframe statistics (stats) and query columns (query), Modin outperforms other libraries on larger datasets exploiting its multithreading power, which even makes ModinD 10 times faster than Polars when running the entire stage on Taxi.
-Overall, Pandas maintains good performance throughout the stage.
+Figure 2.
+Average speedup over Pandas computed for each preparator of the three data preparation pipelines per datasets.
+For each preparator, we report the number of calls in each pipeline and its average impact on the stage execution time.
 (a) Athlete, Read
@@
 (d) Taxi, Read
-Figure 3. Average runtime for reading CSV and Parquet files.
+Figure 3. Average runtime for reading CSV and Parquet files (lower is better).
 (e) Athlete, Write
@@
 (h) Taxi, Write
-Figure 4. Average runtime for writing CSV and Parquet files.
-Data Trasformation (DT)
-In this stage, CuDF emerges as the best performer on Loan (Figure 1(f)) and Patrol (Figure 1(g)), while SparkSQL and Polars register the best performance (also thanks to lazy evaluation) on Athlete (Figure 1(e)) and Taxi (Figure 1(h)), respectively.
-The most time intensive preparators in this stage are categorical encoding (catenc), calculate new column (apply), and group dataframe (group).
-For catenc and group, CuDF always shows the best performance, being at least up to 100 times faster than the second best performer (Polars and ModinR, respectively).
-Once again, Vaex consistently outperforms the other libraries on column-wise operations (e.g., apply), being ${10^{3}}$ times faster than the second best performer CuDF (Figure 2(d)).
-On the other hand, this causes a drastic drop in its performance when moving to operations such as group, join, or pivot.
-As in the previous stage, SparkPD results as the worst performer on several preparators.
-This is mainly due to the latency introduced by translating Pandas operations into Spark execution plans.
-Since SparkSQL does not require this translation, it often achieves significantly better performance.
-When datasets scale up, libraries like ModinD and ModinR, designed for resource-intensive applications, exhibit superior performance due to their optimized handling of large-scale data compared to their performance on smaller datasets.
+Figure 4. Average runtime for writing CSV and Parquet files (lower is better).
+Explorative Data Analysis (EDA)
+For EDA, Polars clearly stands out as the best performer on all datasets (Figures 1(a)-1(d)).
+As depicted in Figure 2, Polars always registers the best performance for isna and outlier, the preparators with the greatest impact on Athlete and Loan, respectively.
+For the former preparator, Polars is up to 10,000 times faster than Pandas and 3 times faster than the second-best performer, i.e., DataTable.
+Both libraries avoid element-wise comparisons by using a special encoding to efficiently track null values: while Polars relies on the validity bitmap used by most types of Arrow arrays in their metadata (Polars, 2024c), DataTable encodes them with sentinel values (H2O, 2024).
+For the latter preparator, outliers are located by filtering the dataframe using the values returned by the quantile function.
+Pandas relies on NumPy exact quantile computation (involving sorting and interpolation) (Pandas, 2024c), while Polars and Spark employ approximate methods that avoid sorting (Apache Spark, 2024b).
+The performance gap between SparkSQL and SparkPD arises from differences in their underlying implementations.
+Although Vaex tends to be very efficient in column-wise operations (e.g, srchptn) and can handle filtering well by internally tracking selected rows without copying data, operations based on percentiles are penalized by the complexity of the required calculations (i.e., determining min/max column values, cumulative sums, and grid interpolation) (Vaex, 2024b).
+For sort, CuDF and Polars are the fastest libraries overall.
+While CuDF leverages on high-performance C++ parallel algorithms from the Thrust library (NVIDIA, 2024b), Polars achieves high performance through its efficient multi-threaded Rust implementation (Polars, 2024e).
+SparkSQL shows remarkable advantages over Pandas as the datasets grow larger, while SparkPD performs similarly to the baseline.
+Modin performs significantly worse than Pandas on Athlete as it requires to apply Pandas implementation of the sorting preparator within each data partition (Modin, 2024d), introducing significant latency to partition and merge results.
+For stats, DataTable outperforms Pandas by up to 50 times by computing statistics either during the creation of the Frame object or efficiently on-demand when needed, while Modin (with both Ray and Dask) is up to 200 times faster than the baseline on Taxi thanks to its multi-threading capabilities.
+Vaex achieves notable performance for query, for which it excels on the Taxi dataset and is only outperformed by Polars, which enables fast querying by leveraging mask operations (Polars, 2024d), on the smaller ones.
+Finally, Pandas generally maintains overall good performance for getcols and dtypes, where PySpark is much slower due to the overhead associated to its inherently distributed architecture, optimized for complex tasks rather than simple ones.
+Data Transformation (DT)
+In this stage, CuDF and Polars emerge as the best performers overall, highlighting the benefits of GPU acceleration and Rust code optimizations, respectively.
+In particular, CuDF outperforms Polars on Loan (Figure 1(f)) and Patrol (Figure 1(g)), while the opposite is true on Athlete (Figure 1(e)).
+As depicted in Figure 2, CuDF always achieves the best results for catenc, while for onehot and group it is the best performer on all datasets but Athlete, where Polars outperforms the alternatives.
+SparkSQL shows excellent results on Patrol (Figure 1(g)): despite not excelling for single preparators (Figure 2(c)), lazy evaluation provides remarkable benefits.
+Moreover, it clearly outperforms SparkPD, which results as the worst performer on several preparators, mainly due to the latency introduced by translating Pandas operations into Spark execution plans.
+Vaex optimizations lead to outstanding performance on column-wise preparators (e.g., calccol and drop), but also to drastic drops moving to operations such as group, join, or pivot.
+Modin, designed for resource-intensive applications, improves its performance as datasets scale up, with ModinR being the best performer for pivot on Taxi.
+DataTable registers remarkable performance on that preparator too, despite not supporting it natively (DataTable, 2024a), but particularly excels for cast, due to direct memory manipulation and in-place casting (DataTable, 2024b).
+On the other hand, libraries relying on the Arrow columnar format (e.g., Polars), with its safety check and abstraction layers, may experience additional overhead (Polars, 2024a).
 Data Cleaning (DC)
-When moving to data cleaning operations, CuDF consistently outperforms other libraries on all datasets, except for Taxi (Figure 1(l)), where Vaex stands out as the best performer (with remarkable results achieved by Polars and Modin with both engines).
-The pipelines designed for the smallest datasets (Athlete and Loan) include a preparator, deduplicate rows (dedup), accounting for 70% of the overall execution time for the stage.
-Pandas register good results on these datasets, especially on Athlete, where only CuDF performs better (Figure 1(i)).
-On Loan, it is possible to notice how Vaex equals or (even significantly) exceeds the performance of CuDF for multiple preparators (Figure 2(b)), but the greater overall impact of the ones where it is outperformed by CuDF (especially the great difference in deduplicating rows) leads to a consistent difference in the overall execution time for this stage (Figure 1(j)).
-Results change when moving to larger datasets (Figures 1(k) and 1(l)).
-CuDF excels on Patrol, where it is over ${10^{3}}$ times faster than Pandas for both considered preparators, i.e., delete empty rows (dropna) and change date & time format (chdate), slightly outperforming ModinR.
-Nevertheless, it struggles on Taxi, where Vaex achieves the best results instead, followed closely by Modin (with both engines) and Polars.
+When it comes to DC, Polars achieves the best results on Athlete (Figure 1(i)), outperforming CuDF (which performs better for all preparators if considered separately, as depicted in Figure 2(a)) thanks to lazy evaluation.
+CuDF excels on Loan (Figure 1(j)) thanks to its performance with dedup, which has by far the highest runtime among the preparators, thus deserving an in-depth analysis in the following.
+CuDF implements dedup natively, by using factorization to identify duplicates (Pandas, 2024b).
+Differently, Polars tracks and manages the presence or absence of unique values using bitmasking (Polars, 2024d), while SparkSQL uses grouping and aggregations to find duplicate rows (Apache Spark, 2024c).
+Finally, Vaex and DataTable lack an official implementation for dedup.
+Results change when moving to larger datasets (Figures 1(k) and 1(l)), with Vaex achieving the best results both on Patrol (followed closely by CuDF and ModinR) and on Taxi (with remarkable performance by Polars and both Modin versions).
+On Patrol, Vaex is the best performer on both preparators (Figure 2(c)): dropna, where it exploits its efficient filtering mechanisms, and chdate, for which it relies on NumPy operations (faster compared to more complex operations found in other libraries) to enhance conversion and manipulation of columns (Vaex, 2024c).
 (a) Athlete
@@
 (d) Taxi
-Figure 5. Average runtime for the execution of the entire data preparation pipelines.
+Figure 5. Average speedup over Pandas for the execution of the entire data preparation pipelines.
 Input/Output (I/O)
 Loading data represents a fundamental step for the success of any data-driven process.
-In particular, the CSV format has established itself as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion (Vitagliano et al., 2023).
-Nowadays, the diffusion of Arrow is pushing towards an increasing adoption of the Parquet format [^25] https://parquet.apache.org, known for its efficient data compression.
-In particular, developers often load Parquet data into memory and convert it to Arrow (Liu et al., 2023), recognized as an ideal in-memory transport layer for such data [^26] https://arrow.apache.org/docs/python/parquet.html.
-Figures 4 and 4 show performance in read and write operations on both file formats, CSV and Parquet.
-DataTable, which maps the file into memory and navigates through it using pointers, records the best overall performance in reading CSV files, but it does not support the Parquet format.
-Polars is the best performer in reading Parquet files due to its capability to load them directly into memory using Arrow.
-When dealing with small datasets like Athlete, Parquet and CSV often exhibit almost the same performance, unlike Polars and SparkSQL (Figure 3(a)).
+In particular, the CSV format is considered as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion (Vitagliano et al., 2023).
+Nowadays, the diffusion of Arrow is pushing towards an increasing adoption of the Parquet format (Apache Parquet, 2024), known for its efficient data compression.
+In particular, developers often load Parquet data into memory and convert it to Arrow (Liu et al., 2023), recognized as an ideal in-memory transport layer for such data (Apache Arrow, 2024b).
+Figures 4 and 4 show performance in reading and writing both file formats, CSV and Parquet.
+When reading small datasets like Athlete, Parquet and CSV often exhibit similar performance, with exceptions such as Polars or SparkSQL (Figure 3(a)).
 In some cases (e.g., Vaex or CuDF), the use of Parquet can even lead to worse performance, due to the need for decompressing the data.
 When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with Polars is over 100 times faster than reading CSV files (Figure 3(d)).
-Write operations exhibit a wide range of performance, mainly due to the need of compressing data before saving.
-Parquet files come with additional metadata and encoding, which may introduce some overhead for the writing process.
-Nevertheless, Parquet performs better than CSV across all datasets as it integrates various efficient compression and encoding approaches, while CSV is plain text and uncompressed (Liu et al., 2023).
-Table 4. Specifications of each machine configuration.
-Laptop Workstation Server # CPUs 8 16 24 # RAM (GB) 16 64 128 Dask (workers - threads) 4-8 4-16 6-24 Ray (workers) 8 16 24
+Overall, CuDF and Vaex appear to be the best performers in reading CSV files.
+Vaex is particularly fast thanks to chunked reading, conversion to optimized formats (e.g., HDF5), and efficient use of memory (Vaex, 2024a).
+CuDF, relying on GPU acceleration and efficient memory management, maintains similar performance for both file formats.
+DataTable, which maps the file into memory and navigates through it using pointers, performs good in reading CSV files, but it does not support Parquet.
+Polars and Vaex excel in reading Parquet files thanks to their capability to load them directly into memory using Arrow.
+Write operations exhibit a wide range of performance.
+Overall, Polars and CuDF are the top performers and Parquet performs better than CSV across all datasets as it integrates efficient compression and encoding approaches, while CSV is plain text and uncompressed (Liu et al., 2023).
+Nevertheless, the overhead introduced by compression and encoding operations can penalize Parquet in some scenarios, as highlighted by CuDF on small datasets (e.g., Athlete and Loan).
+Further, Polars presents a reported issue in writing Parquet files, which determines slower performance and larger file size (Polars, 2024d).
+Finally, Modin (with both engines) demonstrates the best performance on Parquet, particularly for large datasets, by leveraging parallel processing across multiple cores (Modin, 2024a).
 4.2. Impact of Lazy Evaluation
 {mdframed}
-SummaryCuDF significantly outperforms other libraries when running the entire pipeline, while SparkSQL achieves remarkable results on all datasets.
-Lazy evaluation brings benefits mostly on large datasets, improving performance by 30% on average (up to almost 70% in some cases) compared to its eager counterpart.
-Figure 5 reports for each library the runtime for the execution of the entire data preparation pipeline.
+SummaryCuDF generally outperforms other libraries when running the entire pipeline, while SparkSQL and ModinR achieve remarkable results on all datasets.
+Lazy evaluation often brings relevant benefits, improving performance by 20% on average (up to almost 80%) over its eager counterpart.
+Figure 5 reports the average speedup over Pandas achieved by its alternatives for the execution of the entire data preparation pipelines on each dataset.
 For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL, and Polars), we also point out the difference compared to eager evaluation.
-Considering the overall performance on the entire pipeline, CuDF consistently outperforms other libraries by effectively exploiting the power of GPU, with the only exception of the smallest dataset (Figure 5(a)).
-Pandas performs quite well on smaller datasets, but struggles on larger ones due to its eager approach (which requires computing intermediate results) causing higher memory consumption (Sinthong and Carey, 2021).
-Vaex and SparkSQL perform well on Patrol thanks to parallel processing, which allows them to efficiently exploit available CPU cores for handling many string columns.
+Considering the overall performance on the entire pipeline, CuDF generally outperforms other libraries by effectively exploiting the power of GPU, with the only exception of Athlete (Figure 5(a)), where Polars is the best performer.
+PySpark (except for SparkPD on Athlete), ModinR, and Polars show very solid performance overall, as well as Vaex on the largest datasets.
+Pandas performs relatively well on the smallest datasets, but it is outperformed by all alternatives on larger ones, due to its eager approach causing high memory consumption (Sinthong and Carey, 2021).
+Polars generally achieves good results with both kinds of evaluations, as its eager API is highly optimized and in many cases it internally relies on its lazy counterpart before immediately collecting the result.
+Lazy evaluation leverages techniques such as streaming processing, early filtering, and projection pushdown, achieving a performance improvement of up to 25% on Patrol (Figure 5(c)).
 On one hand, it can be noticed how long logical plans can limit the optimizations introduced by lazy evaluation.
-Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement of 30% on Taxi and almost no changes in performance on Patrol.
-On the other hand, libraries like SparkPD leverage Pandas-specific optimizations (e.g., vectorized operations and memory efficiency) that translate into a 70% performance improvement.
-Further, Polars generally demonstrates good performance.
-In many cases, the eager API internally utilizes the lazy API before immediately collecting the result.
-Lazy execution in Polars leverages techniques such as streaming processing, early filtering, and projection pushdown, achieving a performance improvement of up to 15%.
-However, it is important to notice that these optimizations are not always available, hence in some cases lazy operations may perform comparably to their eager counterpart (Figures 5(b) and 5(d)).
-(a) Laptop
-(b) Workstation
-(c) Server
-Figure 6. Average runtime for running the entire pipeline on incremental samples of Taxi.
+Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement of 40% on Taxi and almost no changes in performance on Loan.
+Libraries like SparkPD, leverage to Pandas-specific optimizations (e.g., vectorized operations and efficient memory management) that translate for instance into a 80% performance improvement on Patrol (Figure 5(c)).
 4.3. Scalability
@@
 In this section, we assess how dataframe libraries scale by varying the size of the dataset and the configuration of the underlying machine.
-In particular, we take into account three different machine configurations, whose CPU and RAM are incrementally increased as described in Table 4 to reproduce typical configurations for laptops, workstations, and servers.
-Figure 6 shows the performance of the libraries for running the entire pipeline on incremental samples of Taxi.
-Table 5 outlines instead the minimum configuration required by each library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi.
+In particular, we take into account three different machine configurations, whose CPU and RAM are incrementally increased as described in Table 4 (which also reports the specifications of the Dask and Ray engines used by Modin) to reproduce typical configurations for laptops, workstations, and servers.
+Table 4. Specifications of each machine configuration.
+Laptop Workstation Server # CPUs 8 16 24 # RAM (GB) 16 64 128 Dask (workers - threads) 4-8 4-16 6-24 Ray (workers) 8 16 24
+For this experiment, we selected the most computationally expensive pipeline among the three evaluated, which is the first one [^6] On average, this pipeline is approximately three times more computationally expensive than the others..
 Since CuDF scales according to GPU memory, it is not included in this experiment.
@@
 However, this dependency does not make it a good candidate for dealing with very large datasets on machines with limited GPU resources.
-Overall, SparkSQL clearly stands out as the best performer, being the only library capable of handling both datasets in their entirety on a laptop configuration.
-This capability is due to the use of disk spillover, which offloads data from RAM to disk storage, when the memory limit is reached during computations, ensuring continuous operation without exceeding memory capacity.
-On the other hand, SparkPD suffers instead from overhead due to internal conversions between Pandas and Spark dataframes and increased JVM memory usage, causing out-of-memory errors on the laptop configuration for both Patrol and Taxi.
-DataTable stands out on Patrol as one of the libraries that require the least resources, followed closely by ModinR and Pandas.
-Indeed, it is one of the libraries that consumes the least memory on Patrol; further, it does not require to split the dataframe into multiple pieces, which can lead to errors due to communication and information exchange.
-ModinR demonstrates better scalability across all machine configurations compared to its Dask counterpart.
-The Dask engine uses a centralized scheduler to distribute data across multiple cores, which results in higher memory consumption and leads to out-of-memory (OOM) issues more easily.
-In contrast, Ray can scale far beyond Dask due to its distributed task scheduling scheme, specifically employing a distributed bottom-up scheduling approach.
-This optimizes resource utilization and minimizes overhead by initiating tasks at the lowest level of computation and aggregating results upwards.
-Notably, Modin was originally built to work on top of Ray, making this integration more mature and optimized.
-Despite its excellent performance in other experiments, scalability appears to be a significant weakness for Polars.
+Figure 6 shows the performance of the libraries for running the entire pipeline on incremental samples of Taxi.
+SparkSQL clearly stands out as the best performer, being the only library capable of handling the full pipeline execution on laptop configuration (Figure 6(a)).
+This capability is attributed to its combination of lazy evaluation and disk spillover mechanisms (Apache Spark, 2024e)—optimizing execution plans and offloading data to disk when memory limits are reached during computations.
+SparkPD suffers instead from the overhead due to internal conversions between Pandas and Spark dataframes and increased JVM memory usage, causing out-of-memory (OOM) errors on the laptop configuration.
+While Modin (both Dask and Ray) supports operations on datasets larger than available memory (Modin, 2024a), the efficacy of this feature is constrained by two primary factors.
+Firstly, it requires additional memory for managing internal operations, metadata, and intermediate results.
+Secondly, certain operations (e.g., apply lambda function), necessitate in-memory data access.
+DataTable shares this limitation as well, as a results these two libraries encounter OOM issues when processing datasets larger than 25% of the full dataset.
+Similarly, Vaex experiences OOM errors with datasets exceeding 15% of the full dataset size.
+This is due to the fact that in Vaex the output of some operations (e.g., groupby) is held entirely in memory, potentially exceeding available RAM (Vaex, 2024c).
+Despite Polars excellent performance in other experiments, scalability appears its weakness—it reaches OOM with 4 million and 40 million rows in Figures 6(a) and  6(b), respectively.
 This limitation arises from its in-memory execution model, which requires all data to be loaded into RAM.
@@
 Finally, Pandas confirms its well-known severe limitations about scalability, being the only library not able to complete the pipeline on Taxi even on a server configuration (Figure 6(c)).
-Table 5. Minimum machine configuration for running the entire pipeline on incremental dataset samples. Laptop, Workstation, Server. See Table 4 for the detailed configurations.
+To provide a comprehensive overview of how libraries scale across datasets with different characteristics, Table 5 outlines instead the minimum configuration required by each library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi.
+SparkSQL is the only library able to run the pipelines with the laptop configuration on the entire datasets—thanks to lazy evaluation and disk spillover.
+The second best in scalability is DataTable, which excels on Patrol requiring minimal resources through efficient memory mapping on disk.
+(a) Laptop
+(b) Workstation
+(c) Server
+Figure 6. Average runtime for running the entire pipeline on incremental samples of Taxi.
+ModinR demonstrates better scalability across all machine configurations compared to its Dask counterpart.
+The Dask engine uses a centralized scheduler to distribute data across multiple cores, which results in higher memory consumption and leads to OOM issues more easily (Dask, 2024a).
+In contrast, Ray can scale far beyond Dask due to its distributed task scheduling scheme, specifically employing a distributed bottom-up scheduling approach (Ray, 2024a).
+This optimizes resource utilization and minimizes overhead by initiating tasks at the lowest level of computation and aggregating results upwards.
+Notably, Modin was originally built to work on top of Ray, making this integration more mature and optimized (Modin, 2024a).
+(a) Queries 1 to 11.
+(b) Queries 12 to 22.
+Figure 7. Performance of the dataframe libraries on the TPC-H 10GB queries (lower is better).
+4.4. Performance on the TPC-H Queries
+{mdframed}
+SummaryCuDF consistently achieves the best performance across all queries, while Polars significantly outperforms other CPU-only dataframe libraries.
+TPC-H (TPC-H, 2024) is a decision support benchmark consisting of a suite of business-oriented ad-hoc queries and concurrent data modifications (Pöss and Floyd, 2000).
+It is widely adopted to compare end-to-end database systems (Dreseler et al., 2020), allowing to comprehensively evaluate their performance and efficiency through the execution of complex queries on large volumes of data.
+Consistently with previous related work (Polars, 2024f) exploiting TPC-H queries to assess the performance of a subset of dataframe libraries (see Section 5 for more details), we also adopt this benchmark to provide further support to the outcomes of the comparative analysis presented above.
+In particular, we take as a reference a publicly available translation of the TPC-H queries into Pandas API operations (Bodo, 2024), replicating it for the other libraries, then use Bento for their seamless execution.
+Figure 7 illustrates the performance of each library on the 22 queries of the TPC-H 10GB benchmark.
+We select a scale factor of 10, because it represents the largest dataset that can be processed by the 40 GB RAM of our GPU, which therefore would not be able to handle a scale factor of 100.
+Moreover, we included DuckDB (Raasveldt and Mühleisen, 2019), being a popular in-process SQL analytical database management system.
+DuckDB can execute parallel queries directly on Pandas DataFrames, Parquet/CSV files, or Arrow tables without a separate import step, and can write results back to these formats.
+This interoperability allows seamless integration with other data science libraries.
+We executed the TPC-H queries on DuckDB using the provided implementation in the Polars benchmark (Polars, 2024f).
+Notice that DuckDB does not provide Pandas API and only supports SQL.
+Thus, by expressing the pipelines as SQL queries, it can take advantage of all the well-known query optimization and execution strategies typical of relational database management systems.
+For this reason, we do not compare it with the other libraries throughout the paper, but only report its performance with TPC-H to provide a valuable reference point w.r.t.
+OLAP database management systems.
+Overall, CuDF consistently achieves the best performance by leveraging GPU power.
+Among the CPU-only libraries, Polars is instead the clear winner.
+This confirms what is observed in the data preparation pipelines, for EDA and DT.
+SparkSQL  outperforms Pandas on several queries, registering notable results, especially on q06, where it is even faster than Polars (note that this query selects line items shipped within a year interval, reducing the dataset by 85%).
+By leveraging lazy evaluation, SparkSQL significantly reduces the amount of data processed early in the query execution—this is one of the key reasons why it improves performance in data preparation pipelines by approximately 60%.
+Pandas is never the worst performer, but on most queries, it is among the slowest.
+Modin (which registers the same results with both engines, Dask and Ray) shows diverse performance instead, strongly dependent on the query, confirming the inconsistent performance observed across the different data preparation stages.
+As mentioned above, SparkPD suffers from the additional latency due to the translation of Pandas operations into Spark execution plans, making it one of the worst performers.
+DataTable is one of the worst performers too, since it is slow in grouping operations (as also testified by the execution of group preparator in DT stage and their own benchmark (H2O, 2021)) and its API only supports joins on columns with unique values, requiring therefore to switch to the default Pandas API (e.g., q09).
+Vaex is by far the worst performer overall, as it is significantly slow in grouping operations (as also pointed out in the previous sections) and lacks support for multi-column joins.
+Table 5. Minimum machine configuration for running the entire pipeline on incremental dataset samples.
 Patrol Taxi
@@
 # Rows (M) 0.2 1.2 4.5 6.7 13.5 20.2 27 0.7 3.8 11.5 19.2 38.8 57.7 77
-Pandas LP LP LP LP WS WS WS LP LP WS WS SV SV
-SparkPD WS WS WS WS WS WS WS WS WS WS WS WS WS WS
-SparkSQL LP LP LP LP LP LP LP LP LP LP LP LP LP LP
-ModinD LP LP LP WS SV SV SV LP LP LP WS SV
-ModinR LP LP LP LP WS WS WS LP LP LP LP WS WS WS
-Polars LP LP LP WS WS WS WS LP LP WS WS WS SV SV
-Vaex LP LP LP WS WS WS WS LP LP LP WS WS WS WS
-Datatable LP LP LP LP LP WS WS LP LP LP LP WS WS WS
-4.4. Performance on the TPC-H queries
-{mdframed}
-SummaryCuDF consistently achieves the best performance across all queries, while Polars significantly outperforms other CPU-only dataframe libraries.
-TPC-H [^27] https://www.tpc.org/tpch/default5.asp is a decision support benchmark consisting of a suite of business-oriented ad-hoc queries and concurrent data modifications (Pöss and Floyd, 2000).
-It is widely adopted to compare end-to-end database systems (Dreseler et al., 2020), allowing to comprehensively evaluate their performance and efficiency through the execution of complex queries on large volumes of data.
-Consistently with previous related work [^28] https://pola.rs/posts/benchmarks/ exploiting TPC-H queries to assess the performance of a subset of dataframe libraries (see Section 5 for more details), we also adopt this benchmark to provide further support to the outcomes of the comparative analysis presented above.
-In particular, we take as a reference a publicly available translation of the TPC-H queries into Pandas API operations [^29] https://github.com/Bodo-inc/Bodo-examples, replicating it for the other libraries, then use Bento for their seamless execution.
-Figure 7 illustrates the performance of each library on the 22 queries of the TPC-H 10GB benchmark.
-We select a scale factor of 10, because it represents the largest dataset that can be processed by the 40 GB RAM of our GPU, which therefore would not be able to handle a scale factor of 100.
-Overall, CuDF consistently achieves the best performance by leveraging GPU power.
-Among the CPU-only libraries, Polars is instead the clear winner.
-SparkSQL significantly outperforms Pandas on several queries, registering notable results especially on q06, where it is even faster than Polars (note that this query selects line items shipped within a year interval, reducing the dataset by 85%).
-Pandas is not the worst performer, but on most queries it does not achieve good results.
-Modin (which registers the same results with both engines, Dask and Ray) shows diverse performance instead, strongly dependent on the query.
-As mentioned above, SparkPD suffers from the additional latency due to the translation of Pandas operations into Spark execution plans, making it one of the worst performers.
-DataTable is one of the worst performers too, since it is slow in grouping operations (as also testified by their own benchmark [^30] https://h2oai.github.io/db-benchmark/) and its API only supports joins on columns with unique values, requiring therefore to switch to the default Pandas API (e.g., q09).
-Finally, Vaex is by far the worst performer overall, as it is significantly slow in grouping operations (as also pointed out in the previous sections) and lacks support for multi-column joins.
+Pandas
+SparkPD
+SparkSQL
+ModinD
+ModinR
+Polars
+Vaex
+DataTable
+III Laptop   III Workstation   III Server   OOM
+After CuDF, the second best performers are DuckDB and Polars.
+They excel in different aspects of the benchmark, with neither consistently outperforming the other across all queries—they adopt two different approaches for query optimizations.
 5. Related Work
@@
 The major claims by such analysis are mostly confirmed by our evaluation; nevertheless, despite often providing some useful insights, they only offer a partial and very fragmented knowledge, far from a complete and detailed overview covering all libraries and operations.
-In literature, when a paper introduces a new dataframe library or dataframe-based tool, it usually includes a comparison against some dataframe libraries.
-It is the case for the AFrame data analysis package (Sinthong and Carey, 2019) (later extended to PolyFrame (Sinthong and Carey, 2021)), designed to efficiently manage large-scale dataframes using SQL++ queries.
-AFrame is evaluated against Pandas, Spark, and Pandas on Ray (i.e., ModinR) through a micro-benchmark using the synthetic Wisconsin benchmark dataset (DeWitt, 1993) to assess acceleration and scalability across distributed environments.
-Similarly, Petersohn et al.
+In the literature, only partial comparisons have been performed among libraries, which we list in the following.
+AFrame (Sinthong and Carey, 2019, 2021) reports an evaluation against Pandas, Spark, and ModinR through a micro-benchmark using the synthetic Wisconsin benchmark dataset (DeWitt, 1993) to assess acceleration and scalability across distributed environments.
+Petersohn et al.
 (Petersohn et al., 2021) evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas.
-However, these comparisons share significant limitations, since data preparation is not their main focus and only a relatively small set of libraries and operations is considered.
-Further, dataframe libraries are sometimes compared in data science benchmarks.
-For instance, as previously mentioned in Section 4.4, it is the case for TPC-H (Pöss and Floyd, 2000), which is used by Polars  [^31] https://pola.rs/posts/benchmarks/ to compare against Pandas, PySpark, Dask, Modin, and DuckDB (Raasveldt and Mühleisen, 2019).
-While the outcome of their analysis is consistent with our findings, their evaluation (still performed by selecting a scale factor of 10) only covers the first 7 queries out of the entire batch of 22 covered by our evaluation.
-Similarly, H2O (creator of DataTable) uses a set of queries to asses the scalability of popular database and dataframe-like systems [^32] https://h2oai.github.io/db-benchmark only limited to group by and join operations.
-Indeed, due to their different focus, such relational benchmarking systems notoriously offer very limited insights into real-world data preparation workflows executed on dataframes (Rehman and Elmore, 2022).
-Note that for producing a consistent a analysis we do not take into account tools that are not based on the dataframe data structure, as the aforementioned DuckDB or similar SQL-based solutions.
-For instance, Grizzly (Kläbe and Hagedorn, 2021) aims to address Pandas limitations in scalability by translating dataframes to SQL code, while PyFroid translates from the Pandas API to the DuckDB API to efficiently scale Pandas workloads on a commodity workstation, using top-voted Kaggle notebooks for the evaluation.
-Sanzu (Watson et al., 2017) is designed to evaluate five popular data science frameworks and systems (R, Anaconda [^33] https://www.anaconda.com, Dask, PostgreSQL [^34] https://www.postgresql.org, and PySpark) on data processing and analysis tasks.
-Sanzu is composed of a micro-benchmark, to test basic operations in isolation on synthetic datasets of different sizes, and a macro-benchmark, to evaluate series of operations representing concrete data science scenarios on real-world datasets.
-FuzzyData (Rehman and Elmore, 2022) is a workflow generation system that allows to compare dataframe-based APIs on workflows composed of a small subset of operations.
-Finally, Shanbhag et al.
-(Shanbhag and Chimalakonda, 2023) measure the energy consumption of 21 dataframe processing tasks on Pandas, Vaex, and Dask.
-(a) Queries 1 to 11.
-(b) Queries 12 to 22.
-Figure 7. Performance of the dataframe libraries on the TPC-H 10GB queries.
+Grizzly (Kläbe and Hagedorn, 2021) evaluates against Modin and Pandas, while translating dataframe pipelines to SQL.
+Shanbhag et al.
+(Shanbhag and Chimalakonda, 2023) compare Pandas, Vaex and Dask on a set of single operations to gain insights into their energy consumption.
+Dataframe libraries have been compared also in data science benchmarks.
+For instance, Polars (Polars, 2024f) compares against Pandas, PySpark, Dask, Modin, and DuckDB on TPC-H (Pöss and Floyd, 2000).
+The outcome of their evaluation is consistent with our findings (we adopt the same scale factor for TPC-H), but their evaluation only covers the first 7 queries out of 22.
+Similarly, DataTable (H2O, 2021) uses a set of queries to asses the scalability of popular database and dataframe-like systems only limited to group by and join operations.
+Sanzu (Watson et al., 2017) proposes a benchmark designed to evaluate five popular data science frameworks and systems (R, Anaconda (Anaconda, 2024), Dask, PostgreSQL (PostgreSQL, 2024), and PySpark) on data processing and analysis tasks.
+It comprises both a micro-benchmark for testing basic operations in isolation and a macro-benchmark for evaluating series of operations representing concrete data science scenarios on real-world datasets.
+Similarly, FuzzyData (Rehman and Elmore, 2022) is a workflow generation system that allows to compare dataframe-based APIs on workflows composed of a small subset of operations.
+Other works have explored the problem of optimizing/re-writing dataframe pipelines, reporting comparative evaluations of popular libraries.
+For instance, PyFroid (Emani et al., 2024) translates from the Pandas API to the DuckDB API to efficiently scale Pandas workloads on a commodity workstation, using top-voted Kaggle notebooks for the evaluation.
+Dataprep (Peng et al., 2021) is a task-centric EDA tool that can use different dataframe libraries as back-end engines, including Modin, PySpark, and Dask.
+Dias (Baziotis et al., 2024) is a system for dynamically rewriting Pandas code to optimize performance.
+All of these studies offer valuable insights into dataframe library performance across some real-world scenarios, but they do not provide a comprehensive evaluation on data preparation tasks with all dataframe libraries considered here.
 6. Key Takeaways
 Our evaluation confirmed the well-known limitations of Pandas.
-However, despite the abundance of alternative dataframe libraries designed to overtake it, there is no silver bullet to perform data preparation of tabular data with dataframes on a single machine.
-The right choice usually depends on many factors, such as the size and the features of the dataset at hand, the configuration of the underlying machine, and the kind of operations to carry out.
-We can help data practitioners to decide which library to use by answering the following questions.
+However, despite the abundance of dataframe libraries designed to overtake it, there is no silver bullet to perform the preparation of tabular data with dataframes on a single machine.
+The relative performance of the libraries varies significantly among the variegate datasets and stages of the data preparation pipelines that we considered.
+None of the tools emerges as the clear winner for all considered scenarios.
+Nevertheless, certain characteristics of the dataset and the pipeline can help narrow the user’s choice of the tool.
+Below, we list three questions to assist in this selection.
 Does the dataset fit in the GPU memory?
-If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API.
-However, CuDF lacks of an optimizer; thus, libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations [^35] https://pola-rs.github.io/polars/user-guide/misc/alternatives/.
+If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API [^7] This assumption is based on a single-machine scenario.
+It is important to note that while GPUs can provide substantial performance advantages, comparing GPUs and CPUs requires careful consideration of factors like costs (e.g., pay-per-use fees in a cloud environment) and the growing availability of high-core-count CPUs..
+However, CuDF lacks of an optimizer; thus, libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations (Polars, 2024b)
 Individual stage or complete pipeline?
-If a user is only interested in an individual stage, i.e., input/output (I/O), explorative data analysis (EDA), data transformation (DT), or data cleaning (DC), DataTable stands out as the fastest library for I/O operations, while for other tasks it might not be that mature.
-Further, it shows the least compatibility with Pandas, requiring to implement many preparators manually.
-Polars emerges as the best performer for EDA, while SparkSQL represents a consistent solution for DT and DC, also recording the best performance among all CPU-only libraries when running the entire pipeline.
+If a user is only interested in an individual stage, i.e., input/output (I/O), explorative data analysis (EDA), data transformation (DT), or data cleaning (DC), CuDF and Polars stand out as the fastest libraries for reading and writing operation respectively.
+Polars emerges as the best performer for EDA, while SparkSQL represents a consistent solution for DT.
+Moreover, for DC, Vaex records the best performance among all CPU-only libraries when running the entire pipeline on larger datasets.
 What is the size of the datasets?
 For small datasets (less than 500k rows), all of the libraries show similar average runtimes.
-Even if Polars is often recommended for such a scenario, and indeed registers the best overall performance if memory resources are not limited, Pandas might still be the most reasonable choice, since their performance does not differ substantially—at least, not enough to learn a new API and replace entire parts of the pipeline (Pinner, 2023).
-In the case of datasets of medium size (from 2 to 19 million rows) with dense numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed.
-On the other hand, for sparse datasets with many null values (e.g., Loan), Polars and SparkPD show good performance.
+Even if Polars is often recommended for such a scenario, and indeed registers the best overall performance if memory resources are not limited, Pandas might still be the most reasonable choice, as their performance does not differ substantially—at least, not enough to replace entire parts of the pipeline (Pinner, 2023).
+In the case of datasets of medium size (from 2 to 19 million rows) primarily composed of numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed.
+On the other hand, for datasets with a high proportion of missing values (e.g., Loan, where more than 30% of values are missing), ModinR and PySpark (both version) show good performance.
 Finally, for datasets with many string columns (e.g., 25% Patrol), SparkSQL appears to be the best choice.
@@
 To guarantee comparable results and increase usability, we developed Bento, a general framework for assessing the performance of dataframe libraries on four major data preparation stages: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC).
-We exploited Bento to perform a thorough comparative analysis of the libraries on heterogeneous datasets (varying in size, complexity, and features) previously adopted in literature and publicly available on Kaggle.
-In particular, we relied on the most voted Kaggle notebooks to assess the performance of the libraries on real-world data preparations pipelines validated by a large community.
-Further, we also compare their performance on the popular TPC-H benchmark to support our findings.
-The results of our experiments show that there is no clear winner among the libraries, whose performance is strongly dependent on the size and features of the dataset at hand, on the specifications of the underlying machine, and on the kind of operations to carry out.
-Thus, we provided the readers with useful takeaways to support them in the selection of the right library for the data preparation task at hand.
+We exploited Bento to perform a thorough comparative analysis of the libraries on four heterogeneous datasets (varying in size, complexity, and features) previously adopted in literature and publicly available on Kaggle.
+In particular, we relied on three of the most voted Kaggle notebooks per dataset to assess the performance of the libraries on real-world data preparation pipelines validated by a large community.
+Further, we also evaluate their performance on the popular TPC-H benchmark to support our findings, including DuckDB to assess performance in comparison with an SQL-based system.
+Our experimental evaluation shows that one size does not fit all when it comes to dataframe libraries for data preparation.
+However, we have distilled key takeaways that data practitioners can use as starting points for selecting the appropriate library for their task at hand.
 As future work, we plan to compare the libraries in a distributed environment and study the possible adoption of machine learning to suggest optimal library combinations based on datasets, tasks, hardware, and previous executions.
+Further, we plan to explore how the libraries perform in combination with tools that optimize and rewrite entire dataframe-based data preparation pipelines, such as Dataprep (Peng et al., 2021) and Dias (Baziotis et al., 2024).
 Acknowledgements
-We extend our appreciation to the master’s students of the 2021/2022 Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions in testing the libraries.
+We extend our appreciation to the master’s students of the 2021-2022 Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions in testing the libraries.
 We warmly thank Leonardo S.p.A., co-funding the PhD program of Angelo Mozzillo.